### v1

# Counting points on character varieties

Abstract.
We count points on the character varieties associated with punctured surfaces and regular semisimple generic conjugacy classes in reductive groups. We find that the number of points are palindromic polynomials. This suggests a $P=W$ conjecture for these varieties. We also count points on the corresponding additive character varieties and find that the number of points are also polynomials, which we conjecture have non-negative coefficients. These polynomials can be considered as the reductive analogues of the Kac polynomials of comet-shaped quivers.
Key words and phrases:
Character varieties, representation varieties, additive character varieties, $E$-polynomials, generic conjugacy classes
2010 Mathematics Subject Classification:
14M35, 14D23, 11G25
1. Introduction
1.1. Overview
Character varieties of surface groups play a central role in diverse areas of mathematics and physics, such as gauge theory [AtiyahBott83, HitchinProceedingofLMS87], mirror symmetry [HauselThaddeusInventiones2003], non-abelian Hodge theory [Simpson92, Simpson94i], and the geometric Langlands program [BeilinsonDrinfeld, BenZviNadler]. Consequently, the topology and geometry of these varieties have been the focus of active research for decades. In their groundbreaking works [HRV08, HLRV11], Hausel, Letellier, and Rodriguez-Villegas counted points on $\mathrm{GL}_{n}$-character varieties, leading to deep conjectures about the geometry of these varieties. Many of these conjectures have since been proven, greatly elucidating the geometry of character varieties. Notable achievements include the determination of the Betti numbers [SchiffmannAnnals2016, MellitInventiones21, Mellit20], the establishment of curious Poincaré duality and curious Hard Lefschetz [MellitDecomposition], and the recent resolution of the $P=W$ conjecture, cf. [HoskinPW].
In [HLRV11], the authors also considered the additive version of character varieties, where the natural commutation relation is replaced by its Lie algebraic analogue. According to the Riemann–Hilbert correspondence, points on character varieties correspond to flat connections on vector bundles. When $g=0$, the points on the additive character variety correspond to flat connections on the trivial bundle. Thus, when $g=0$, the additive character variety is a dense open subvariety of the usual character variety. This direct relationship is lost in higher genus; nevertheless, it was conjectured in [HLRV11] that the cohomology of the additive character variety captures the pure part of the cohomology of the multiplicative character variety in all genus. As evidence, the authors proved that additive character varieties are pure. This result relied on the beautiful bridge built by Crawley-Boevey between additive character varieties and quiver varieties associated to comet-shaped quivers [CBIndecomposable]. Under this correspondence, the number of points on the additive character variety matches the Kac polynomials of the corresponding quiver. Utilising these ideas, Hausel, Letellier, and Rodriguez-Villegas proved Kac’s conjecture for arbitrary quivers [HLVKacAnnals2013].
Despite significant advances in understanding the geometry of character varieties associated with $\mathrm{GL}_{n}$ (and more generally, groups of type $A$), there is limited understanding of character varieties associated with arbitrary reductive groups $G$. Understanding these general cases is crucial for applications in fields such as Langlands duality and mirror symmetry, where the Langlands dual group plays a central role. In this paper, we count points on both multiplicative and additive character varieties associated with punctured surface groups and semisimple conjugacy classes in arbitrary reductive groups. Our work builds upon and generalises the contributions of Cambò [Cambo17], who studied $\mathrm{Sp}_{2n}$-character varieties; [BK22], who investigated character stacks associated with compact surface groups; and [KNP23], who examined character varieties in presence of regular semisimple and regular unipotent conjugacy classes at the punctures.
One of our main results is that the number of points of multiplicative character varieties over finite fields are palindromic polynomials. This hints at a $P=W$ conjecture, a curious Poincaré duality, and a curious Hard Lefschetz property for these varieties. [^1] To make further progress in this direction, one needs a reductive analogue of Markman’s result [MarkmanGenerators2001] that the tautological classes (along with those from the underlying surface) generate the cohomology. Another important contribution of this paper is an explicit polynomial formula for the number of points of additive character varieties. The resulting polynomials can be considered as the reductive analogues of the Kac polynomials of comet-shaped quivers. Computer calculations suggest that these polynomials have non-negative coefficients [^2] See [Bailey] for the computer code written using CHEVIE.; however, lack of direct connection with quiver varieties makes proving this a nontrivial task. Finally, one of the main contributions of [HRV08, HLRV11] was to relate the number of points of $\mathrm{GL}_{n}$-character varieties with Macdonald polynomials and use this to provide a conjectural formula for the mixed Hodge polynomials of these varieties. [^3] As far as we know, these conjectures (as well as the purity conjecture mentioned above) are still open. Formulating analogous conjectures in the reductive setting is an interesting open problem.
1.1.1. Structure of the text
In the rest of the introduction, we formulate our main results. In §2, we gather some facts about reductive groups and root systems. In §3, we discuss basic results about the geometry of generic character varieties. The main computation regarding points on character varieties appears in §4 while its additive analogue appears in §5. Finally, in §6, we give explicit examples of counting polynomials for groups of low rank and discuss some of the consequences of these formulas.
1.1.2. Acknowledgement
We would like to thank Ivan Cheltsov, Anand Deopurkar, Jack Hall, Ian Le, Emmanuel Letellier, Paul Levy, Jean Michel, and Behrouz Taji for helpful conversations. MK was supported by an Australian Research Council (ARC) discovery grant. GH was supported by the National Research Foundation of Korea (NRF). BW and SG were supported by Australian postgraduate scholarships. We would like to thank the Sydney Mathematical Research Institute for facilitating a visit by Emmanuel Letellier, giving further impetus to this project. The material of §4 (resp. §5) forms a part of the Master’s thesis of BW (resp. SG).
1.2. Generic character varieties
Let $g$ and $n$ be integers satisfying $g\geq 0$ and $n\geq 1$. Let $\Gamma=\Gamma_{g,n}$ be the fundamental group of an orientable surface with genus $g$ and $n$ punctures:
$\Gamma=\frac{\langle a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n}\rangle}{[a_{1},b_{1}]\cdots[a_{g},b_{g}]c_{1}\cdots c_{n}}.$
Let $G$ be a connected split reductive group over $\mathbb{F}_{q}$ with split maximal torus $T$ and centre $Z$. Let $C=(C_{1},\ldots,C_{n})$ be a tuple of conjugacy classes of $G$. The representation variety associated to $(\Gamma,G,C)$ is defined by
$R:=\bigg{\{}(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})\in G^{2g}\times\prod_{i=1}^{n}C_{i}\ \bigg{|}\ \prod_{i=1}^{g}[a_{i},b_{i}]\prod_{i=1}^{n}c_{i}=1\bigg{\}}.$
This is an affine scheme of finite type over $\mathbb{F}_{q}$ with an action of $G$ by conjugation. The associated character variety is
${X}:=R/\!\!/G.$
We also have the corresponding character stack
$[{X}]:=[R/(G/Z)].$
Assumption 1.
We will be working under the following assumptions:
  1. (1)
    $\prod_{i=1}^{n}C_{i}\subseteq[G,G]$ (otherwise, $R$ would be empty).
  2. (2)
    Each conjugacy class $C_{i}$ contains an element $S_{i}\in T(\mathbb{F}_{q})$ which is strongly regular; i.e., $C_{G}(S_{i})=T$.
  3. (3)
    The tuple $C=(C_{1},\ldots,C_{n})$ is generic in the sense of §3.1.
We also assume $2g+n\geq 3$, otherwise the generic assumption implies that $X$ is empty. As explained in §3.2, the generic assumption ensures that the action of $G/Z$ on $R$ has finite étale stabilisers and that $R$ is smooth and equidimensional. Thus, ${X}$ is also equidimensional of pure dimension
$\dim({X})=(2g-2+n)\dim(G)+2\dim(Z)-n.\dim(T).$
We assume $Z=Z(G)$ is connected, the ground characteristic is very good for $G$ and $q$ is large enough (see Convention 25). Then, we have:
Theorem 2.
$|X(\mathbb{F}_{q})|$ is a polynomial in $q$ given explicitly in §4.7. This is the $E$-polynomial [^4] See, e.g., [LetellierVillegas] for recollections on $E$-polynomials of varieties over $\mathbb{F}_{q}$. of $X$ and depends only on $(G,g,n)$; i.e., it is independent of the conjugacy classes $C_{i}$ (provided they satisfy Assumption 1). Moreover, this polynomial is palindromic.
The starting point of the proof is the observation that since $X$ is the coarse moduli space of the Deligne-Mumford stack $[X]$, we have $|X(\mathbb{F}_{q})|=|[X](\mathbb{F}_{q})|$. An application of Lang’s theorem then implies
$|[X](\mathbb{F}_{q})|=\frac{|R(\mathbb{F}_{q})||Z(\mathbb{F}_{q})|}{|G(\mathbb{F}_{q})|}.$
Next, the Frobenius Mass Formula allows one to write $|R(\mathbb{F}_{q})|$ in terms of irreducible complex representations of the finite group $G(\mathbb{F}_{q})$. We then use results of Deligne and Lusztig to analyse this formula and prove the theorem; see §4 for details.
Remark 3.
  1. (i)
    We expect that the theorem holds without assuming that $Z(G)$ is connected; however, the expression for $|X(\mathbb{F}_{q})|$ becomes more complicated due to disconnected centralisers in the dual group.
  2. (ii)
    Our analysis shows that $|[{X}](\mathbb{F}_{q})|$ is a polynomial in $q$ even if we drop the generic assumption. We expect that the regularity assumption is also not required for this. We emphasise, however, that without the generic assumption, the variety $X$ is quite different from the stack $[X]$.
  3. (iii)
    The leading coefficient of the $E$-polynomial equals the number of connected components of ${X}$. It is easy to analyse this polynomial to conclude that $|\pi_{0}({X})|=|\pi_{0}(Z(\check{G}))|$, where $\check{G}$ is the Langlands dual group of $G$; see [Whitbread24, §7.5] for details.
  4. (iv)
    The value at $1$ of the $E$-polynomial gives the Euler characteristic of $X$. This can be used to show that if $g>1$, or $g=1$ and $Z$ is nontrivial, then the Euler characteristic of ${X}$ is 00. Expressions for the Euler characteristic when $g=0$ or $g=1$ and $Z$ is trivial are given in §6.1. For instance, when $G=G_{2}$ and $g=0$, we find
    $\chi(X)=2^{2n-7}3^{n-3}(n-1)(n-2)(207n^{2}-621n+350).$
In lieu of the above theorem, it is reasonable to expect:
Conjecture 4.
The analogue of ${X}$ over the complex numbers is Hodge–Tate, and satisfies the curious Poincaré duality and curious hard Lefschetz. Moreover, there is a version of the $P=W$ conjecture for ${X}$.
We expect that the above theorem and conjecture hold without the regularity assumption. However, the generic assumption and semisimplicity are crucial. When $G=\mathrm{GL}_{n}$, Mellit constructed a cell decomposition for ${X}$, thus providing a conceptual understanding of the fact that ${X}$ is polynomial count and proving in addition that it is Hodge–Tate and satisfies curious Hard Lefschetz and curious Poincaré duality [Mellit20]. We expect that ${X}$ has an analogous cell decomposition for arbitrary reductive group $G$.
1.3. Additive analogue
Let $\mathfrak{g}$ be the Lie algebra of $G$ and ${\mathcal{O}=(\mathcal{O}_{1},\ldots,\mathcal{O}_{n})}$ a tuple of $G$-orbits on $\mathfrak{g}$. Define the additive representation variety by
${A}:=\bigg{\{}(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})\in\mathfrak{g}^{2g}\times\prod_{i=1}^{n}\mathcal{O}_{i}\ \bigg{|}\ \sum_{i=1}^{g}[a_{i},b_{i}]+\sum_{i=1}^{n}c_{i}=0\bigg{\}}.$
We have the adjoint action of $G$ on ${A}$. We call the resulting GIT quotient
${Y}:={A}/\!\!/G$
the additive character variety associated to $(\Gamma,G,\mathcal{O})$. We will be working under the obvious additive analogues of Assumption 1. Then the action of $G/Z$ on ${A}$ has finite étale stabilisers, ${A}$ is smooth and equidimensional, and ${Y}$ is also equidimensional of the same dimension as ${X}$. As above, we assume that $Z$ is connected, the ground characteristic is very good and $q$ is large enough (see Convention 26). Then we have:
Theorem 5.
$|Y(\mathbb{F}_{q})|$ is a polynomial in $q$ given explicitly in §5.6. This is the $E$-polynomial of $Y$ and depends only on $(G,g,n)$; i.e., it is independent of the orbits $\mathcal{O}_{i}$ (provided they satisfy the additive analogue of Assumption 1). Moreover, this polynomial is monic; thus, $X$ is connected.
The starting point of the proof is an additive analogue of the Frobenius formula which gives an expression for $|{Y}(\mathbb{F}_{q})|$ in terms of the Fourier transforms of $G$-invariant orbits on $\mathfrak{g}$. We then use a result of Kazhdan and Letellier to evaluate these Fourier transforms and compute $|Y(\mathbb{F}_{q})|$; see §5 for details. Using the formula for $|Y(\mathbb{F}_{q})|$, we can compute the Euler characteristic $Y$. For instance, for $G=G_{2}$, we find that $\chi(Y)$ equals
$\frac{1}{6}(6-2^{n}+19\cdot 3^{n+1}+5^{n})+395\cdot 12^{n-2}+5g\cdot 6^{n}+6^{n-2}(48-65\cdot 2^{n})n+5\cdot 3^{n}4^{n-2}(16g^{2}+8g(n-3)+n^{2}).$
When $G=\mathrm{GL}_{n}$, the additive character variety $Y$ is isomorphic to a quiver variety and $|Y(\mathbb{F}_{q})|$ is a polynomial with non-negative coefficients [HLRV11]. Computer calculations suggest that this fact generalises to the reductive setting:
Conjecture 6.
The variety ${Y}$ is pure; thus, $|{Y}(\mathbb{F}_{q})|$ has non-negative coefficients.
As noted in the overview section, when $G=\mathrm{GL}_{n}$, $|{Y}(\mathbb{F}_{q})|$ is the Kac polynomial of a certain comet shaped quiver. The coefficients of Kac polynomials are expected to be the graded dimensions of the associated graded Borcherd algebra [Schiffmann18, §3]. It would be interesting to find a similar interpretation for the coefficients of $|{Y}(\mathbb{F}_{q})|$ for general reductive groups.
2. Root systems and reductive groups
In this section, we recall basic facts and notations about root systems and reductive groups. The material here is standard, though perhaps the facts about pseudo-Levi subgroups and quasi-isolated semisimple elements are not as widely known.
2.1. Root systems
[…]
2.1.1. Subsystems
A root subsystem of $\Phi$ is a subset $\Psi\subseteq\Phi$ which is itself a root system. This is equivalent to the requirement that for all $\alpha,\beta\in\Psi$, $s_{\alpha}(\beta)\in\Psi$. A subsystem $\Psi\subseteq\Phi$ is closed if
$\alpha,\beta\in\Psi\quad\textrm{and}\quad\alpha+\beta\in\Phi\quad\implies\quad\alpha+\beta\in\Psi.$
Given a subset $S\subseteq\Phi$, let $\langle S\rangle_{\mathbb{Z}}$ denote the subgroup of $V$ generated by $S$. The subsystem of $\Phi$ generated by $S$ is defined by
$\langle S\rangle:=\langle S\rangle_{\mathbb{Z}}\cap\Phi.$
Note that $\langle S\rangle$ is a closed subsystem of $\Phi$.
2.1.2. Dual root systems
For each root $\alpha\in\Phi$, define the coroot $\check{\alpha}:=\frac{2}{(\alpha,\alpha)}\alpha\in V$. The set of coroots forms the dual root system $\check{\Phi}\subset V$. If $\Psi$ is a subsystem of $\Phi$, then $\check{\Psi}$ is a subsystem of $\check{\Phi}$. However, if $\Psi$ is closed in $\Phi$, then $\check{\Psi}$ need not be closed in $\check{\Phi}$. For instance, $C_{n}\times C_{n}$ is a closed subsystem of $C_{2n}$ but if $n>1$, then $B_{n}\times B_{n}$ is not closed in $B_{2n}$. Indeed, it is easy to see that $B_{n}\times B_{n}$ does not arise in the Borel–de Siebenthal algorithm applied to $B_{2n}$.
2.1.3. Levi subsystems
A Levi subsystem [^5] In the literature, these are usually called parabolic subsystems, but we find our terminology more useful for comparing with pseudo-Levi and endoscopy subsystems defined below. of $\Phi$ is a subsystem of the form $\Phi\cap E$ where $E\subseteq V$ is a subspace. A Levi subsystem of $\Phi$ is closed in $\Phi$. Moreover, every Levi subsystem of $\Phi$ is of the form $w.\langle S\rangle$, where $S$ is a subset of $\Delta$ and $w\in W$. If $w=1$, then $\langle S\rangle$ is called a standard Levi subsystem. Note that $\Psi$ is a Levi subsystem of $\Phi$ if and only if $\check{\Psi}$ is a Levi subsystem of $\check{\Phi}$.
2.1.4. Isolated subsystems
A subset $S\subseteq\Phi$ is called isolated in $\Phi$ if $S$ is not contained in a proper Levi subsystem of $\Phi$. One can check that the following are equivalent:
  1. (1)
    $S$ is isolated in $\Phi$.
  2. (2)
    $\check{S}$ is isolated in $\check{\Phi}$.
  3. (3)
    $S$ does not lie in a proper subspace of $V$.
  4. (4)
    $\langle S\rangle_{\mathbb{Z}}\otimes\mathbb{R}=V$.
  5. (5)
    The root systems $\langle S\rangle$ and $\Phi$ have the same rank.
  6. (6)
    $\bigcap_{\alpha\in S}\mathrm{ker}(\alpha)=0$.
2.1.5. Pseudo-Levi subsystems
Assume $\Phi$ is irreducible. Let $\theta$ be the highest root of $\Phi$, and $\widetilde{\Delta}:=\Delta\sqcup\{-\theta\}$. A pseudo-Levi subsystem is a subsystem of $\Phi$ of the form $w.\langle S\rangle$, where $w\in W$ and $S$ is a subset of $\widetilde{\Delta}$. If $w=1$, then $\langle S\rangle$ is called a standard pseudo-Levi subsystem. More generally, if $\Phi=\Phi_{1}\sqcup\cdots\sqcup\Phi_{r}$ with $\Phi_{i}$ irreducible, then a subsystem $\Psi\subseteq\Phi$ is said to be a pseudo-Levi if $\Psi_{i}:=\Psi\cap\Phi_{i}$ is a pseudo-Levi inside $\Phi_{i}$ for all $i$. It is clear that every Levi is a pseudo-Levi. The converse is true in type $A$ but false otherwise. For instance, $C_{2}$ has a pseudo-Levi $A_{1}\times A_{1}$ which is not a Levi.
2.1.6. Isolated pseudo-Levis
Let $\Psi$ be a pseudo-Levi subsystem of an irreducible root system $\Phi$. One can check that the following are equivalent:
  1. (1)
    $\Psi$ is isolated in $\Phi$.
  2. (2)
    $\Psi=w.\langle S\rangle$ where $S\subset\tilde{\Delta}$ has the same size as $\Delta$.
In other words, up to $W$-conjugation, isolated pseudo-Levis are those which are obtained by removing just a single element of $\widetilde{\Delta}$ (equivalently, a single node from the extended Dynkin diagram). We leave it to the reader to generalise these statements to non-irreducible root systems. Figure 1 lists the isolated pseudo-Levi subsystems of irreducible root systems.
Type of $\Phi$ Isolated pseudo-Levi subsystems of $\Phi$
$A_{n}$ $(n\geq 1)$ $A_{n}$ only
$B_{n}$ $(n\geq 3)$ $B_{n}$,  $D_{n}$,  $D_{n-1}\times A_{1}^{\prime}$,  $D_{r}\times B_{n-r}$ ($22\leq r\leq n-$ 2)
$C_{n}$ $(n\geq 2)$ $C_{n}$,  $C_{r}\times C_{n-r}$ ($1\leq r\leq\frac{n}{2}$)
$D_{n}$ $(n\geq 4)$ $D_{n}$,  $D_{r}\times D_{n-r}$ ($2\leq r\leq\frac{n}{2}$)
$G_{2}$ $G_{2}$,  $A_{2}$,  $A_{1}\times A_{1}^{\prime}$
$F_{4}$ $F_{4}$,  $B_{4}$,  $C_{3}\times A_{1}$,  $A_{3}\times A_{1}^{\prime}$,  $A_{2}\times A_{2}^{\prime}$
$E_{6}$ $E_{6}$,  $A_{5}\times A_{1}$,  $A_{2}\times A_{2}\times A_{2}$
$E_{7}$ $E_{7}$,  $A_{7}$,  $D_{6}\times A_{1}$,  $A_{5}\times A_{2}$,  $A_{3}\times A_{3}\times A_{1}$
$E_{8}$ $E_{8}$,  $D_{8}$,  $A_{8}$,  $E_{7}\times A_{1}$,  $A_{7}\times A_{1}$, $E_{6}\times A_{2}$,  $D_{5}\times A_{3}$, $A_{5}\times A_{2}\times A_{1}$,  $A_{4}\times A_{4}$
Figure 1. Isolated pseudo-Levi subsystems.
2.1.7. Endoscopy subsystems
Let $\check{\Psi}$ be a pseudo-Levi subsystem of $\check{\Phi}$. Then the dual root system $\Psi$ is called an endoscopy subsystem of $\Phi$. For instance, $B_{n}\times B_{n}$ is an endoscopy subsystem of $B_{2n}$. Note that this subsystem is not closed.
2.1.8. Very good primes
We say the prime $p$ is good for $\Phi$ if $p$ does not divide the coefficients of the highest root of $\Phi$. Otherwise $p$ is bad for $G$. Bad characteristics are $p=2$ if the root system is of type $B_{n}$, $C_{n}$ or $D_{n}$, $p=2,3$ in type $G_{2}$, $F_{4}$, $E_{6}$, $E_{7}$ and $p=2,3,5$ in type $E_{8}$. We say $p$ is very good for $\Phi$ if $p$ is good for $\Phi$ and $p$ does not divide $n+1$ for every irreducible component of type $A_{n}$ of $\Phi$. Note that $p$ is (very) good for $\Phi$ if and only if it is (very) good for $\check{\Phi}$.
2.2. Reductive groups
Let $G$ be a connected reductive group over an algebraically closed field $k$. Let $T$ be a maximal torus of $G$, $B$ a Borel containing $T$, $U$ the unipotent radical of $B$, $W$ the Weyl group, $(X,\Phi,\check{X},\check{\Phi})$ the root datum, $\langle\Phi\rangle_{\mathbb{Z}}$ root lattice, and $\langle\check{\Phi}\rangle_{\mathbb{Z}}$ the coroot lattice. Let $\Delta\subseteq\Phi$ (resp. $\check{\Delta}\subseteq\check{\Phi}$) denote the base determined by $B$ and $V:=X\otimes\mathbb{R}$.
2.2.1. Dual group
Let $\check{T}:=\mathrm{Spec}\,k[X]$ be the dual torus and $\check{G}$ the Langlands dual group of $G$ over $k$. In other words, $\check{G}$ is the connected reductive group over $k$ with maximal torus $\check{T}$ and root datum $(\check{X},\check{\Phi},X,\Phi)$.
2.2.2. Levi subgroups
A Levi subgroup of $G$ is defined to be the centraliser $L=C_{G}(S)$ of a torus $S\subseteq G$. If $S\subseteq T$ (in other words, if $L$ contains $T$), then $L$ is said to be a standard Levi subgroup. In this case, the root system of $L$ is a Levi subsystem of $\Phi$. Every Levi subsystem arises in this manner.
If $x\in\mathfrak{g}$ is a semisimple element, then the connected centraliser $C_{G}(x)^{\circ}$ is a Levi subgroup of $G$. Moreover, every Levi subgroup of $G$ arises in this manner. Note that if the characteristic of $k$ is very good for $G$, then $C_{G}(x)$ is connected [Steinberg75, Theorem 3.14].
2.2.3. Parabolic subgroups
A connected subgroup $P\subseteq G$ is called parabolic if the quotient $G/P$ is proper. Every parabolic subgroup has a Levi decomposition $P=L\ltimes N$, where $N$ is the maximal closed connected normal unipotent subgroup of $P$ and $L$ is a Levi subgroup of $G$. Every Levi subgroup of $G$ arises in this manner. In what follows, we will be using the following facts:
  1. (1)
    If $s\in P$ is semisimple, then $s$ is $P$-conjugate to an element of $L$. Indeed, $s$ lies in some maximal torus $T^{\prime}$ of $P$ and since all maximal tori of $P$ are $P$-conjugate, we can choose a $P$-conjugate of $T^{\prime}$ which lies in $L$.
  2. (2)
    If $l\in L$, then $l\in[L,L]$ if and only if $l\in[P,P]$. This follows from the fact that $[P,P]=[L,L]\ltimes N$.
2.2.4. Pseudo-Levi subgroups
Let $x\in G$ be a semisimple element. The connected centraliser $G_{x}^{\circ}=C_{G}(x)^{\circ}$ is called a pseudo-Levi subgroup of $G$. The dual group is called an endoscopy group for $G$. If $x\in T$, then $G_{x}^{\circ}$ is called a standard pseudo-Levi subgroup of $G$ and its dual is called a standard endoscopy group for $\check{G}$.
2.2.5.
If $x\in T$, then the root system of $G_{x}^{\circ}$ is given by
${\Phi}_{x}:=\{{\alpha}\in{\Phi}\,|\,{\alpha}(x)=1\}.$
This is a pseudo-Levi subsystem of $\Phi$. Moreover, every pseudo-Levi subsystem arises via the centraliser of some element of $T$, cf. [Deriziotis]. By a theorem of Steinberg, $\pi_{0}(G_{x})$ is isomorphic to a subgroup of $\pi_{1}([G,G])$; thus, if $[G,G]$ is simply connected, then $G_{x}$ is connected [Steinberg75, §2].
2.2.6. Quasi-isolated elements
A semisimple element $x\in G$ is called quasi-isolated if $G_{x}$ is not in a proper Levi subgroup of $G$. It is called isolated if $G_{x}^{\circ}$ is not in any proper Levi subgroup of $G$. Thus, $x\in T$ is isolated if and only if $\Phi_{x}$ is an isolated (pseudo-Levi) subsystem of $\Phi$. If $x$ is quasi-isolated, then $G_{x}$ is an irreducible subgroup of $G$ (i.e., it does not belong to a proper parabolic). By Schur’s lemma (cf. [Sikora, Proposition 15]) $Z(G_{x})/Z(G)$ is finite.
2.2.7. Pretty good primes
Following [Herpel], we say a prime $p$ is pretty good for $G$ if for every subset $A\subseteq\Phi$, the finitely generated abelian groups $X/\mathbb{Z}A$ and $\check{X}/\mathbb{Z}\check{A}$ have no $p$-torsion.
Theorem 7 ([Herpel]).
  1. (i)
    If $p$ is very good for $G$, then it is pretty good for $G$.
  2. (ii)
    Suppose $\tilde{G}$ is a possibly disconnected reductive group with connected component $G$, the prime $p=\mathrm{char}(k)$ is pretty good for $G$, and $\pi_{0}(\widetilde{G})$ is étale. Then the centraliser $C_{\tilde{G}}(H)$ of every closed subgroup scheme $H\subseteq\tilde{G}$ is smooth.
Corollary 8.
If $p=\mathrm{char}(k)$ is very good for $G$, then for every $x\in T$, the centre $Z(G_{x})$ is smooth.
Proof.
Indeed, $p$ being pretty good for $G$ implies that it is pretty good for $G_{x}^{\circ}$ and that $G_{x}=C_{G}(x)$ is smooth; thus, $\pi_{0}(G_{x})$ is étale. By the previous theorem, $Z(G_{x})=C_{G_{x}}(G_{x})$ is smooth. ∎
3. Generic character varieties
Let $G$ be a connected reductive group over an algebraically closed field $k$. In this section, we define the notion generic tuples of conjugacy classes of $G$ and study the corresponding character varieties. Most of this material appears in [HLRV11] and [Boalch14], but our exposition is slightly different and more suited for our purposes.
3.1. Generic classes
Let $C_{1},\ldots,C_{n}$ be conjugacy classes of $G$.
Definition 9.
The tuple $C=(C_{1},\ldots,C_{n})$ is called generic if whenever there exists a proper parabolic subgroup $P\subset G$ and $x_{i}\in P\cap C_{i}$, for $i=1,\ldots,n$, then $\prod_{i=1}^{n}x_{i}\notin[P,P]$.
3.1.1. Maximal parabolics
Clearly it is enough to check the condition on maximal proper parabolic subgroups. If $G=\mathrm{GL}_{n}$, this is equivalent to requiring that if a proper non-zero subspace $V\subset k^{n}$ is preserved by $x_{i}\in C_{i}$, $i=1,\ldots,n$, then $\prod_{i=1}^{n}\mathrm{det}(x_{i}|_{V})\neq 1$; thus, we recover the notion of generic in [HLRV11, §2].
3.1.2. Standard parabolics
It is enough to check the above condition for standard (proper) parabolic subgroups. Here, standard parabolic subgroups are parabolic subgroups containing a chosen Borel subgroup $B\subset G$. Indeed, suppose $P$ is an arbitrary parabolic and choose $g\in G$ such that the conjugate $g.P$ is standard. Then $x_{i}\in P\cap C_{i}$ if and only if $g.x_{i}\in g.P\cap C_{i}$ and $\prod_{i=1}^{n}x_{i}\in[P,P]$ if and only if $g.\prod_{i=1}^{n}x_{i}=\prod_{i=1}^{n}g.x_{i}\in g.[P,P]=[g.P,g.P]$.
3.1.3. Semisimple part
For each conjugacy class $X$, write $X^{s}$ for the corresponding semisimple class, obtained by taking the semisimple parts of the elements of $X$. Then one can show that $C=(C_{1},\ldots,C_{n})$ is generic if and only if $C^{s}=(C_{1}^{s},\ldots,C_{n}^{s})$ is generic. This is the reductive analogue of the fact that, in $\mathrm{GL}_{n}$, being generic depends only on the eigenvalues.
3.1.4. Levis instead of parabolics
For semisimple conjugacy classes, one can check genericity using Levi subgroups instead of parabolics. More precisely, it follows from discussions of §2.2.3 that a tuple of semisimple conjugacy classes $C=(C_{1},\ldots,C_{n})$ is generic if and only if whenever there exists a proper standard Levi subgroup $L\subset G$ and $s_{i}\in L\cap C_{i}$, for $i=1,\ldots,n$, then $\prod_{i=1}^{n}s_{i}\notin[L,L]$.
3.1.5. Generic elements
We now explain that deciding whether a tuple of (semisimple) conjugacy classes is generic involves checking finitely many conditions.
Definition 10.
A tuple $S=(S_{1},\ldots,S_{n})\in T^{n}$ is said to be generic if for every $w=(w_{1},\ldots,w_{n})\in W^{n}$ and every proper standard Levi subgroup $L\subset G$, we have $\prod_{i=1}^{n}w_{i}.S_{i}\notin T\cap[L,L]$.
Let $C_{i}:=G.S_{i}$ be the conjugacy class corresponding to $S_{i}$.
Lemma 11.
The element $S=(S_{1},\ldots,S_{n})\in T^{n}$ is generic if and only if the tuple of conjugacy classes $C=(C_{1},\ldots,C_{n})$ is generic.
Proof.
It is clear that if $C$ is generic then so is $S$. For the converse, let $x=(x_{1},\ldots,x_{n})\in C_{1}\times\ldots\times C_{n}$. Let $L\subset G$ be a proper standard Levi subgroup containing $x_{1},\ldots,x_{n}$. We need to show $\prod_{i=1}^{n}x_{i}\notin[L,L]$. We have $x_{i}=l_{i}.t_{i}$ for some $l_{i}\in L$ and $t_{i}\in T$. Now $t_{i}$ and $S_{i}$ are elements of $T$ which are $G$-conjugate; thus, they must actually be conjugate under $W$; i.e., there exists $w_{i}\in W$ such that $t_{i}=w_{i}.S_{i}$. By assumption, $t_{1}\cdots t_{n}\notin[L,L]$ which implies that $x_{1}\cdots x_{n}\notin[L,L]$. Thus, $C$ is generic. ∎
3.1.6. Existence
The above lemma shows that generic conjugacy classes exist. Indeed, if $L$ is a proper standard levi subgroup, then $T\cap[L,L]$ is a proper closed subvariety of $T$. (At the level of Lie algebras, $\mathfrak{t}=\mathfrak{t}\cap[\mathfrak{l},\mathfrak{l}]\oplus Z(\mathfrak{l})$ for $\mathfrak{t}=\mathrm{Lie}(T)$ and the centre $Z(\mathfrak{l})$ is nontrivial if $\mathfrak{l}$ is a proper Levi subalgebra of $\mathfrak{g}$.) Since there are only finitely many standard Levis $L$ and finitely many $w$’s, we conclude that the generic locus of $T^{n}$ is an open dense subvariety.
3.1.7.
Let $\mathbf{T}_{n}\subseteq T^{n}$ denote the subvariety consisting of those tuples $(S_{1},\ldots,S_{n})\in T^{n}$ satisfying $\prod_{i=1}^{n}S_{i}\in[G,G]$. Then the above argument shows that the generic locus of $\mathbf{T}_{n}$ is open dense. Since the (strongly) regular locus is also open dense in $T^{n}$, it follows that conjugacy classes satisfying Assumption 1 exist.
3.2. Generic character varieties
In this section, we assume that $p=\mathrm{char}(k)$ is a very good prime for $G$ and $Z=Z(G)$ is connected. Let $C=(C_{1},\ldots,C_{n})$ be a generic tuple of conjugacy classes of $G$ satisfying $\prod_{i=1}^{n}C_{i}\subseteq[G,G]$. Let $M:G^{2g}\times C_{1}\times\ldots\times C_{n}\rightarrow[G,G]$ be the map
$M(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n}):=[a_{1},b_{1}]\cdots[a_{g},b_{g}]c_{1}\cdots c_{n}.$
The representation variety associated to $(\Gamma_{g,n},G,C)$ is defined by $R:=M^{-1}(1)$. This is an affine scheme of finite type over $k$. We assume that $R$ is non-empty (this is guaranteed if $g>0$).
Theorem 12.
Every element of $R$ is irreducible and stable and the action of $G/Z$ on $R$ has finite étale stabilisers. Moreover, $R$ is smooth and equidimensional with
$\dim(R)=2g\dim(G)+\sum_{i=1}^{n}\dim(C_{i})-\dim([G,G]).$
Proof.
Let ${r=(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})\in R}$ and suppose $P$ is a parabolic subgroup of $G$ containing all $a_{i},b_{i},$ and $c_{i}$. Then, $[P,P]$ contains $[a_{1},b_{1}]\cdots[a_{g},b_{g}]$; thus, $[P,P]$ also contains $\prod_{i=1}^{n}c_{i}$. As $C$ is generic, it follows that $P=G$. Thus, $r$ is irreducible and Schur’s lemma (cf. [Sikora, §4]) implies that $\mathrm{Stab}_{G/Z}(r)$ is finite. Note that this group is the same as the stabiliser of $\langle a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n}\rangle\subseteq G$; thus, by Theorem 7, it is étale.
Stability follows from the fact that whenever a connected reductive group acts on an affine scheme of finite type with finite stabilisers, then every point is stable, cf. [Mumford, Proposition 0.8]. (Alternatively, one can use the Hilbert–Mumford criterion as in [Boalch14].) The statements about $R$ follow from the fact that $1$ is a regular value for $M$, cf. [KNP23, §2] for details. ∎
3.2.1.
Next, let ${X}:=R/\!\!/G$ and $[{X}]:=[R/(G/Z)]$ denote, respectively, the generic character variety and character stack associated to $(\Gamma_{g,n},G,C)$. The above theorem implies:
Corollary 13.
$[{X}]$ is a smooth Deligne-Mumford stack with coarse moduli space ${X}$.
3.3. Frobenius Mass Formula
We now work over $k=\overline{\mathbb{F}_{q}}$ and assume that $G$ is equipped with an untwisted Frobenius automorphism $F:G\rightarrow G$. We also assume that the conjugacy classes $C=(C_{1},\ldots,C_{n})$ are $F$-stable. Then the action of the Frobenius $F$ on $G$ and $C_{i}$’s, gives an action of $F$ on $R$, ${X}$ and $[{X}]$. We can therefore speak about the Frobenius fixed points $R^{F}$, ${X}^{F}$, and $[{X}]^{F}$. The first two are finite sets while the last one is a finite groupoid.
3.3.1. Points of the representation variety
By definition,
$R^{F}=\Big{\{}(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})\in(G^{F})^{2g}\times\prod_{i=1}^{n}C_{i}^{F}\,|\,M(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})=1\Big{\}}.$
Henceforth, we assume that the stabiliser in $G$ of each element in $C_{i}$ is connected. It then follows that $C_{i}^{F}$ is a unique $G^{F}$-conjugacy class, cf. [GM20, §2.6.1]. Thus, it makes sense to evaluate irreducible complex characters of $G^{F}$ on $C_{i}^{F}$. The following result goes back to the work of Frobenius; see e.g. [HLRV11, §3] for a proof:
Theorem 14 (Frobenius Mass Formula).
${|R^{F}|}=|G^{F}|\sum_{\chi\in\mathrm{Irr}(G^{F})}\left(\frac{|G^{F}|}{\chi(1)}\right)^{2g-2}\prod_{i=1}^{n}\frac{\chi(C_{i}^{F})|C_{i}^{F}|}{\chi(1)}.$
3.3.2. Points of the character stack
An application of Lang’s theorem implies that
$|[{X}]^{F}|=\frac{|R^{F}|}{|(G/Z)^{F}|},$
cf. [Behrend91, Lemma 3.5.6]. Thus, the Frobenius Mass Formula also provides an expression for the number of points of the character stack.
3.3.3. Points of character variety
We now assume that the tuple $(C_{1},\ldots,C_{n})$ is generic. We saw above that ${X}$ is a coarse moduli space for the smooth Deligne-Mumford stack $[{X}]$. This implies that $[{X}]$ and $X$ have the same number of points over finite fields, cf. [Behrend91, §2] or [BogaartEdixhoven]. Thus, we have
$|{X}^{F}|=|[{X}]^{F}|=|Z^{F}|\sum_{\chi\in\mathrm{Irr}(G^{F})}\left(\frac{|G^{F}|}{\chi(1)}\right)^{2g-2}\prod_{i=1}^{n}\frac{\chi(C_{i}^{F})|C_{i}^{F}|}{\chi(1)}.$
3.4. Additive character varieties
The above considerations have additive analogues which we leave to the reader to formulate. The only nontrivial modification is the additive analogue of the Frobenius Mass Formula, which we now discuss.
3.4.1.
Let $\mathcal{O}_{1},\ldots,\mathcal{O}_{n}$ be $F$-stable $G$-orbits in $\mathfrak{g}$. Let $A$ denote the corresponding additive representation variety. Then
${A}^{F}=\Big{\{}(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})\in(\mathfrak{g}^{F})^{2g}\times\prod_{i=1}^{n}\mathcal{O}_{i}^{F}\,|\,m(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})=0\Big{\}},$
where $m:\mathfrak{g}^{2g}\times\mathcal{O}_{1}\times\ldots\times\mathcal{O}_{n}\rightarrow[\mathfrak{g},\mathfrak{g}]$ is defined by
$m(a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n})=\sum_{i=1}^{g}[a_{i},b_{i}]+\sum_{i=1}^{n}c_{i}.$
3.4.2.
Let $1_{\mathcal{O}_{i}}^{G}$ denote the characteristic function of the $G^{F}$-orbit $\mathcal{O}_{i}^{F}$. Let $\mathcal{F}$ denote the Fourier transform on the space of $G^{F}$-invariant functions on $\mathfrak{g}^{F}$ (see §5.1 for the definition). For each $x\in\mathfrak{g}$, let $\mathfrak{g}_{x}$ denote its centraliser in $\mathfrak{g}$. Assume that the stabiliser in $G$ of every element of $\mathcal{O}_{i}$ is connected. Thus, $\mathcal{O}_{i}^{F}\subset\mathfrak{g}^{F}$ is a unique $G^{F}$-orbit. Then we have (cf. [HLRV11, §3]):
Theorem 15 (Additive Frobenius Mass Formula).
${|{A}^{F}|}=|\mathfrak{g}^{F}|^{g-1}\sum_{x\in\mathfrak{g}^{F}}|\mathfrak{g}_{x}^{F}|^{g}\prod_{i=1}^{n}\mathcal{F}(1_{\mathcal{O}_{i}}^{G})(x).$
Corollary 16.
If $\mathcal{O}=(\mathcal{O}_{1},\ldots,\mathcal{O}_{n})$ is generic, then
$|Y^{F}|=|[Y]^{F}|=\frac{|{A}^{F}||Z^{F}|}{|G^{F}|}=\frac{|Z^{F}||\mathfrak{g}^{F}|^{g-1}}{|G^{F}|}\sum_{x\in\mathfrak{g}^{F}}|\mathfrak{g}_{x}^{F}|^{g}\prod_{i=1}^{n}\mathcal{F}(1_{\mathcal{O}_{i}}^{G})(x).$
4. Counting points on character varieties
In this section, we count points on the character varieties considered in §1.2. We treat the case of a once-punctured surface first and mention the modifications required in the multi-punctured case at the end.
Let $G$ be a connected split reductive group over $k=\overline{\mathbb{F}_{q}}$ equipped with an untwisted Frobenius $F$. We assume $Z(G)$ is connected and $\mathrm{char}(k)$ is very good for $G$. Let $T$ be a maximal split $F$-stable torus of $G$. Let $S\in T^{F}$ be a strongly regular generic element and $C=G.S$ the corresponding conjugacy class in $G$. Let $X$ denote the character variety associated to the once-punctured Riemann surface and the conjugacy class $C$. By the discussions of the previous section, we have
(1) $|X^{F}|=\frac{|R^{F}|}{|(G/Z)^{F}|}=\frac{|Z^{F}|}{|T^{F}|}\sum_{\chi\in\mathrm{Irr}(G^{F})}\left(\frac{|G^{F}|}{\chi(1)}\right)^{2g-1}\chi(S).$
We now use a crucial result of Deligne and Lusztig to rewrite the above sum in a more tractable form.
4.1. A theorem of Deligne and Lusztig
Given a finite abelian group $A$, write $A^{\vee}$ for the Pontryagin dual $\mathrm{Hom}(A,\mathbb{C}^{\times})$. Recall that a principal series representation is an irreducible constituent of
$R_{T}^{G}\,\theta:=\mathrm{Ind}_{B^{F}}^{G^{F}}\,\theta,\qquad\theta\in(T^{F})^{\vee}.$
Theorem 17 (Corollary 7.6 of [DL76]).
Let $\chi\in\mathrm{Irr}(G^{F})$ be an arbitrary character and $S\in T^{F}$ a strongly regular element. Then
$\chi(S)=\sum_{\theta\in(T^{F})^{\vee}}\langle\chi,R_{T}^{G}\,\theta\rangle\,\theta(S).$
Corollary 18.
If $S\in T^{F}$ is a strongly regular element and $\chi(S)\neq 0$, then $\chi$ is a principal series representation.
The above corollary implies that only principal series representations contribute to the sum in the expression (1) for $|X^{F}|$. Note that one does not need Deligne–Lusztig theory to formulate the above theorem and corollary. However, the only proof we know of uses the full power of Deligne–Lusztig theory.
4.1.1.
Recall that the induced representations $R_{T}^{G}\,\theta$ and $R_{T}^{G}\,\theta^{\prime}$ have a common constituent if and only if $\theta$ and $\theta^{\prime}$ are $W$-conjugate, cf. [DL76, Corollary 6.3]. Thus, we can associate to a principal series representation $\chi\in\mathrm{Irr}(G^{F})$ a character $\theta\in(T^{F})^{\vee}$, well-defined up to $W$-conjugacy.
4.1.2. Interlude on Pontryagin duality
Let $\mu_{\infty,p^{\prime}}(\mathbb{C})$ denote the set of roots of unity in $\mathbb{C}^{\times}$ whose order is prime to $p$. We choose, once and for all, isomorphisms
(2) $k^{\times}\simeq(\mathbb{Q}/\mathbb{Z})_{p^{\prime}}\qquad\textrm{and}\qquad(\mathbb{Q}/\mathbb{Z})_{p^{\prime}}\simeq\mu_{\infty,p^{\prime}}(\mathbb{C}).$
As noted in [DL76, §5], this induces an isomorphism
(3) $(T^{F})^{\vee}\simeq\check{T}^{F}$
4.1.3.
Using the above identification, we can think of $\theta\in(T^{F})^{\vee}$ as an element of $\check{T}^{F}\subseteq\check{G}^{F}$. Let $\check{G}_{\theta}$ denote the centraliser of $\theta$ in $\check{G}$. The fact that $Z(G)$ is connected implies that $\check{G}$ has simply connected derived subgroup. By Steinberg’s theorem, the stabiliser $\check{G}_{\theta}$ is a connected standard pseudo-Levi subgroup of $\check{G}$. The Weyl group of $\check{G}_{\theta}$ equals
$W_{\theta}=\{w\in W\,|\,w.\theta=\theta\}.$
The Langlands dual group $G_{\theta}:=(\check{G}_{\theta})^{\check{}}$ is a standard endoscopy group for $G$, containing $T$, with Weyl group $W(G_{\theta})=W_{\theta}$.
4.1.4.
Since $G$ is assumed to have connected centre, the constituents of $R_{T}^{G}\,\theta$ are in canonical bijection with $\mathrm{Irr}(W_{\theta})$, cf. [Kilmoyer78, Corollary 4.20]. Let $\chi_{\theta,\rho}\in R_{T}^{G}\,\theta$ be the irreducible character corresponding to $\rho\in\mathrm{Irr}(W_{\theta})$. Then, the double centraliser theorem (cf. [Etingof11, Theorem 5.18.1]) gives
$\langle\chi_{\theta,\rho},R_{T}^{G}\,\theta\rangle=\dim(\rho).$
4.1.5.
Consider the set of pairs $(L,\rho)$, where $L$ is a standard endoscopy group for $G$ and $\rho$ is an irreducible character of $W(L)$. Note $W$ acts on this set by conjugation. Let $\mathcal{T}(G)$ denote the set of $W$-orbits. We call the elements of $\mathcal{T}(G)$ $G$-types and denote them by $\tau=[L,\rho]$. [^6] Here, for ease of notation, we write $[L,\rho]$ for the orbit $[(L,\rho)]$. Note that $\mathcal{T}(G)$ is a finite set and depends only on the root system $\Phi$ of $G$; in particular, it is independent of the ground field.
By the above discussions, we can associate a type to every principal series representation. [^7] This is analogous to the types used in [HRV08, HLRV11, Cambo17]; see [Whitbread24, §4] for a detailed comparison between these notions. Using Lusztig’s Jordan decomposition, one can generalise the notion of types to all irreducible $G^{F}$-characters. We shall not need this more general version in the present paper. In other words, we have a map
(4) Principal series representations of $\textrm{Principal series representations of $G^{F}$}\rightarrow\mathcal{T}(G).$
4.1.6. Unipotent characters
As noted above, irreducible constituents of $R_{T}^{L}(1)$ are in canonical bijection with irreducible representations of $W(L)$. Let $\tilde{\rho}$ denote the unipotent (principal series) character of $L^{F}$ corresponding to $\rho\in\mathrm{Irr}(W(L))$.
Definition 19.
We define the $q$-mass of $\tau=[L,\rho]\in\mathcal{T}(G)$ by
$m_{\tau}(q):=q^{|\Phi^{+}(G)|-|\Phi^{+}(L)|}\frac{|L^{F}|}{\tilde{\rho}(1)}.$
Here, $\Phi^{+}(L)$ denotes the set of positive roots of $L$. Note that $\tilde{\rho}(1)$ is a polynomial in $q$ which divides $|L^{F}|$, cf. [GM20, Remark 2.3.27]. Thus, $m_{\tau}(q)$ is a polynomial in $q$.
4.1.7.
The relevance of $m_{\tau}(q)$ emerges in the following proposition which can be found in, e.g. [GM20, Corollary 2.6.6]:
Proposition 20.
Suppose $\chi$ is a principal series character of $G^{F}$ of type $\tau=[L,\rho]$. Then
$\frac{|G^{F}|}{\chi(1)}=m_{\tau}(q).$
4.1.8.
We can now rewrite the Frobenius Mass Formula (1) in terms of types. For each type $\tau\in\mathcal{T}(G)$, let $\mathrm{Irr}(G^{F})_{\tau}$ denote the set of irreducible principal series characters of type $\tau$. Then, we have
$|{X}^{F}|=\frac{|Z^{F}|}{|T^{F}|}\sum_{\tau\in\mathcal{T}(G)}m_{\tau}(q)^{2g-1}\sum_{\chi\in\mathrm{Irr}(G^{F})_{\tau}}\chi(S).$
4.1.9.
For ease of notation, define
$\boxed{S_{\tau}(q):=\displaystyle\sum_{\chi\in\mathrm{Irr}(G^{F})_{\tau}}\chi(S).}$
Then, we find
$|{X}^{F}|=\frac{|Z^{F}|}{|T^{F}|}\sum_{\tau\in\mathcal{T}(G)}m_{\tau}(q)^{2g-1}S_{\tau}(q).$
Thus, computing $|{X}^{F}|$ is reduced to computing the character sums $S_{\tau}(q)$. [^8] In obtaining the above formula for $|{X}^{F}|$, we used the assumption that $S$ is strongly regular, but not the assumption that $S$ is generic.
4.2. Evaluating the character sums $S_{\tau}$
In this subsection, we rewrite the expression for $S_{\tau}$ in terms of sums of characters of $T^{F}$, thus turning the problem of computing $S_{\tau}$ into a “commutative” problem.
Suppose $\tau=[L,\rho]\in\mathcal{T}(G)$ where $L$ is a standard endoscopy group for $G$ with Weyl group $W(L)$ and $\rho$ is an irreducible character of $W(L)$. Let $[L]$ denote the $W$-orbit of $L$ in $G$ and $[W(L)]$ the $W$-orbit of $W(L)$ in $W$.
4.2.1.
We start by elucidating the nature of $\mathrm{Irr}(G^{F})_{\tau}$. It consists of characters $\chi_{\theta,\rho}$ as $\theta$ runs over elements of $(T^{F})^{\vee}$ satisfying $[G_{\theta}]=[L]$. Equivalently,
$\mathrm{Irr}(G^{F})_{\tau}=\{\chi_{\theta,\rho}\,|\,\theta\in(T^{F})^{\vee},\,[W_{\theta}]=[W(L)]\}.$
We refer the reader to §4.1.4 for the definition of $\chi_{\theta,\rho}$.
4.2.2.
The above discussion implies
$S_{\tau}(q)=\displaystyle\sum_{\chi\in\mathrm{Irr}(G^{F})_{\tau}}\chi(S)=\frac{|W(L)|}{|W|}\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ [W_{\theta}]=[W(L)]\end{subarray}}\chi_{\theta,\rho}(S)=\frac{|W(L)|}{|N_{W}(W(L))|}\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ W_{\theta}=W(L)\end{subarray}}\chi_{\theta,\rho}(S).$
The second equality follows from the fact that $\chi_{\theta,\rho}$ depends only on the $W$-orbit of $\theta$.
4.2.3.
Recall that $[L]$ denotes the orbit of $L$ under the action of $W$ on subgroups of $G$. According to [Carter72, Lemma 34], we have ${|W/N_{W}(W(L))|}=|[L]|$. Thus, we obtain
$S_{\tau}(q)=\frac{|W(L)||[L]|}{|W|}\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ W_{\theta}=W(L)\end{subarray}}\chi_{\theta,\rho}(S).$
Remark 21.
In [Carter72, Proposition 28], one finds the following formula for $|[L]|$. Let $W^{\perp}(L)$ be the Weyl group generated by the roots orthogonal to $\Phi(L)$. Let $A_{W}(L)$ denote the group of symmetries of the Dynkin diagram of $L$ induced by elements of $W$. Then
$|[L]|=\frac{|W|}{|W(L)||W^{\perp}(L)||A_{W}(L)|}.$
4.2.4.
Theorem 17 gives the following expression for $\chi_{\theta,\rho}$
$\chi_{\theta,\rho}(S)=\sum_{\theta^{\prime}\in\check{T}^{F}}\langle\chi_{\theta,\rho},R_{T}^{G}\,\theta^{\prime}\rangle\,\theta^{\prime}(S).$
Note that a summand is zero unless $\theta^{\prime}$ is in the same $W$-orbit as $\theta$, in which case it equals $\dim(\rho)\,\theta^{\prime}(S)$; thus,
$\chi_{\theta,\rho}(S)=\dim(\rho)\sum_{w\in W/W_{\theta}}(w.\theta)(S)=\frac{\dim(\rho)}{|W_{\theta}|}\sum_{w\in W}\theta(w.S).$
Note that the last sum depends only on the $W$-orbit of $\theta$.
4.2.5.
We now plug in the explicit formula for $\chi_{\theta,\rho}$ into the formula for $S_{\tau}$ to obtain
$S_{\tau}(q)=\frac{|W(L)||[L]|}{|W|}\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ W_{\theta}=W(L)\end{subarray}}\frac{\dim(\rho)}{|W_{\theta}|}\sum_{w\in W}\theta(w.S)=\frac{\dim(\rho)|[L]|}{|W|}\sum_{w\in W}\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ W_{\theta}=W(L)\end{subarray}}\theta(w.S).$
4.2.6.
To alleviate notation, let
$\boxed{\alpha_{L,S}(q):=\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ W_{\theta}=W(L)\end{subarray}}\theta(S).}$
Thus, we have shown
$S_{\tau}(q)=\frac{\dim(\rho)|[L]|}{|W|}\sum_{w\in W}\alpha_{L,w.S}(q).$
Thus, the problem of computing $S_{\tau}$ reduces to evaluating the character sum $\alpha_{L,S}$. To compute the latter, we first consider an auxiliary sum $\alpha_{L,S}^{\supseteq}$ which is easier to handle.
4.3. The character sum $\alpha_{L,S}^{\supseteq}$
Given a standard endoscopy group $L$ for $G$ and $S\in T^{F}$, define
$\alpha_{L,S}^{\supseteq}(q):=\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ W_{\theta}\supseteq W(L)\end{subarray}}\theta(S)=\sum_{\theta\in(\check{T}^{F})^{W(L)}}\theta(S).$
Lemma 22.
We have
$\alpha^{\supseteq}_{L,S}(q)=\begin{cases}|(\check{T}^{F})^{W(L)}|&\textrm{if $S\in[L^{F},L^{F}]$}\\ 0&\textrm{otherwise}.\end{cases}$
Proof.
This is a straightforward application of the fact that the Pontryagin dual of the embedding of finite abelian groups $(\check{T}^{F})^{W(L)}\hookrightarrow\check{T}^{F}$ is the canonical quotient map
$T^{F}\twoheadrightarrow\frac{T^{F}}{T^{F}\cap[L^{F},L^{F}]}.$
See [KNP23, §5.2] for details. ∎
4.3.1.
Now we use the fact $S\in[G,G]$ is assumed to be generic. This implies that $S\in[L,L]$ if and only if the root system $\Phi(L)$ is isolated in $\Phi=\Phi(G)$; i.e., $\Phi(L)$ is not contained in any proper Levi subsystem of $\Phi$. Now there are finitely many isolated standard endoscopic groups for $G$. Thus, replacing $F$ by $F^{m}$ if necessary, we may assume that $S\in[L^{F},L^{F}]$ for all isolated standard endoscopy groups. Hence, we obtain:
Corollary 23.
Replacing the Frobenius $F$ by a finite power if necessary, we have:
$\alpha^{\supseteq}_{L,S}(q)=\begin{cases}|(\check{T}^{F})^{W(L)}|&\textrm{if $L$ is isolated}\\ 0&\textrm{otherwise}.\end{cases}$
4.3.2. Description of $\check{T}^{W(L)}$
It is proved in [KNP23, §5] that, under the assumption that $G$ has connected centre, we have:
$\check{T}^{W(L)}\simeq Z(\check{L}).$
Thus,
$|(\check{T}^{F})^{W(L)}|=|Z(\check{L})^{F}|=|\pi_{0}(Z(\check{L}))^{F}|\times|(Z(\check{L})^{\circ})^{F}|.$
By Corollary 8, the prime $p=\mathrm{char}(k)$ does not divide $|\pi_{0}(Z(\check{L}))|$. Therefore, replacing $F$ by $F^{m}$ for some positive integer $m$ if necessary, we may assume that $q\equiv 1\mod|\pi_{0}(Z(\check{L}))|$. This implies that $\pi_{0}(Z(\check{L}))^{F}=\pi_{0}(Z(\check{L}))$. Thus, we obtain:
Corollary 24.
Replacing $F$ by a finite power if necessary, we have
$\alpha^{\supseteq}_{L,S}(q)=\begin{cases}|\pi_{0}(Z(\check{L}))|\times|(Z(\check{L})^{\circ})^{F}|&\textrm{if $L$ is isolated}\\ 0&\textrm{otherwise}.\end{cases}$
Convention 25.
Henceforth, we assume $q$ is large enough so that the above equality holds; i.e., $q$ is large enough so that for every isolated standard endoscopy $L$ of $G$, we have $S\in[L^{F},L^{F}]$ and $q\equiv 1\mod|\pi_{0}(Z(\check{L}))|$.
4.4. Evaluating $\alpha_{L,S}$
We now obtain an expression for $\alpha_{L,S}$ using Möbius inversion.
4.4.1.
Let $\mu$ denote the Möbius function on the (finite) poset of endoscopy subsystems of $\Phi$, ordered by inclusion. Möbius inversion implies
$\alpha_{L,S}=\sum_{L^{\prime}}\mu(L,L^{\prime})\alpha^{\supseteq}_{L^{\prime},S},$
Here, the sum is over standard endoscopy groups $L^{\prime}$ whose root system satisfies $\Phi(L^{\prime})\supseteq\Phi(L)$ and we have abused notation and written $\mu(L,L^{\prime})$ for $\mu(\Phi(L),\Phi(L^{\prime}))$.
4.4.2.
We have seen that $\alpha_{L^{\prime},S}^{\supseteq}$ is 00 unless $L^{\prime}$ is isolated, in which case
$Z(\check{L^{\prime}})^{\circ}=Z(\check{G})^{\circ}\simeq Z(G)^{\circ}=Z(G),$
since we assume that $G$ has connected centre, and
$\alpha_{L^{\prime},S}^{\supseteq}=|\pi_{0}(Z(\check{L^{\prime}}))|\times|Z(G)^{F}|.$
4.4.3.
To simplify notation, let
$\nu(L):=\sum_{L^{\prime}}|\pi_{0}(Z(\check{L^{\prime}}))|\,\mu(L,L^{\prime}).$
Here, the sum runs over isolated standard endoscopy groups $L^{\prime}$ whose root systems contain $\Phi(L)$. Thus, $\nu(L)$ is an integer depending only on the root data of $L$ and $G$; in particular, it is independent of the ground field $k$ (of very good characteristic). In fact, $\nu(L)$ depends only on the $W$-orbit of $L$. In view of the above discussions, we conclude
(5) $\boxed{\,\alpha_{L,S}=|Z^{F}|\,\nu(L).\,}$
Observe that $\alpha_{L,S}$ is independent of $S$ (provided $S$ is generic).
4.5. Conclusion of the point count for the once-punctured case
We have seen that
$|{X}^{F}|=\frac{|Z^{F}|}{|T^{F}|}\sum_{\tau=[L,\rho]\in\mathcal{T}(G)}S_{\tau}(q)\,m_{\tau}(q)^{2g-1}.$
Moreover,
$S_{\tau}(q)=\frac{\dim(\rho)\,|[L]|}{|W|}\sum_{w\in W}\alpha_{L,w.S}(q)=\dim(\rho)\,|[L]|\,\nu(L)\,|Z^{F}|.$
Thus, we conclude that $|X^{F}|$ is a rational function in $q$ and that $|X^{F^{m}}|$ is obtained by evaluating the same rational function at $q^{m}$. It follows that this rational function is actually a polynomial. Thus, ${X}$ is polynomial count, in the sense of [LetellierVillegas], and our expression for $|{X}^{F}|$ gives the $E$-polynomial of $X$.
4.6. Palindromic property
Our goal is to explain that the polynomial $|{X}^{F}|$ is palindromic. Recall (cf. [DM20, §7.2]) that the Alvis–Curtis duality is (up to a sign) an involution
$\mathfrak{D}:\mathrm{Irr}(G^{F})\rightarrow\mathrm{Irr}(G^{F}).$
To compute the dimension of the irreducible character $\mathfrak{D}(\chi)$, we need to replace $q$ by $q^{-1}$ in the polynomial $d_{\chi}$ encoding the dimension of $\chi$. More precisely,
$\dim\mathfrak{D}(\chi)=q^{|\Phi^{+}|}d_{\chi}(q^{-1}).$
4.6.1.
It is known that $\mathfrak{D}$ restricts to an involution on the set of irreducible constituents of $R_{T}^{G}\,\theta$; namely, it sends the irreducible constituent corresponding to $\rho\in\mathrm{Irr}(W_{\theta})$ to the constituent corresponding to $\epsilon\rho$, where $\epsilon$ is the sign character of $W_{\theta}$. In particular, we see that $\mathfrak{D}$ sends an irreducible character of type $[L,\rho]$ to one of type $[L,\epsilon\rho]$. Using these facts, it is easy to show that $|{X}^{F}|$ is palindromic. For details, see [Whitbread24, §7.7].
4.7. Multi-punctured case
Let ${X}$ be a character variety associated to an $n$-punctured surface group satisfying Assumption 1. Then a similar analysis as in the once-punctured case gives
$|{X}^{F}|=\frac{|Z^{F}|}{|T^{F}|^{n}}\sum_{\tau\in\mathcal{T}(G)}S_{\tau}(q)\,m_{\tau}(q)^{2g-2+n},$
where
$S_{\tau}(q):=\sum_{\chi\in\mathrm{Irr}(G^{F})_{\tau}}\prod_{i=1}^{n}\chi(S_{i})=\frac{\dim(\rho)^{n}\,|W|^{n-1}\,|[L]|\,\nu(L)}{|W(L)|^{n-1}}|Z^{F}|.$
For details, see [Whitbread24, §6.2].
4.7.1.
For the benefit of the reader, we remind what the notation means. Here, $Z$ is the centre and $T$ is a split maximal torus of $G$, $\mathcal{T}(G)$ is the set of $G$-types consisting of $W$-orbits of pairs $(L,\rho)$, where $L$ is a standard endoscopy group for $G$ and $\rho$ is an irreducible character of $W(L)$. The integer $|[L]|$ is the size of the $W$-orbit of $L$, and the integer $\nu(L)$ is defined in §4.4 using the Möbius function on the poset of endoscopy subsystems of $\Phi$. Finally,
$m_{\tau}(q):=q^{|\Phi^{+}(G)|-|\Phi^{+}(L)|}\frac{|L^{F}|}{\dim(\tilde{\rho})}\in\mathbb{Z}[q],$
where $\tilde{\rho}$ is the unipotent character of $L^{F}$ corresponding to $\rho$.
5. Counting points on additive character varieties
In this section, we count points on the additive character variety introduced in §1.3. We assume first that we have a single puncture and mention the modifications necessary for the multi-punctured case at the end of this section.
5.0.1.
We continue using the notation of the previous section and let $\mathfrak{g}=\mathrm{Lie}({G})$ and $\mathfrak{t}=\mathrm{Lie}(T)$. Let $H\in\mathfrak{t}^{F}$ be a generic regular element and $A$ (resp. $Y$) the additive representation variety (resp. additive character variety) associated to the orbit of $H$ and the once-punctured surface group. By the discussions of §3.4, we have
(6) $|{Y}^{F}|=\frac{|Z^{F}||\mathfrak{g}^{F}|^{g-1}}{|G|^{F}}\sum_{x\in\mathfrak{g}^{F}}|\mathfrak{g}_{x}^{F}|^{g}\mathcal{F}(1_{H}^{G})(x),$
To proceed further, we need to recall some results on the Fourier transforms of invariant functions on $\mathfrak{g}^{F}$.
5.1. Recollections on invariant functions
Fix a nontrivial additive character $\psi:\mathbb{F}_{q}\rightarrow\mathbb{C}^{\times}$ and a non-degenerate $G$-invariant symmetric bilinear form $\kappa:\mathfrak{g}^{F}\times\mathfrak{g}^{F}\rightarrow\mathbb{F}_{q}$. Since we have assumed that the ground characteristic is very good for $G$, such an invariant form exists, cf. [Letellier05, Proposition 2.5.12]. Note that the restriction $\kappa|_{\mathfrak{t}}$ of $\kappa$ to $\mathfrak{t}^{F}$ is also non-degenerate.
5.1.1. Green function
The Green function $Q_{T}^{G}$ is the function from unipotent elements of $G^{F}$ to $\mathbb{Z}$ defined by
$Q_{T}^{G}(u):=\Big{(}\mathrm{Ind}_{B^{F}}^{G^{F}}1\Big{)}(u)=|\mathcal{B}_{u}^{F}|.$
Here, $\mathcal{B}_{u}\subseteq G/B$ denotes the Springer fibre associated to the unipotent element $u\in G^{F}$. Clearly, $Q_{T}^{G}(u)$ depends only on the $G$-orbit of $u$.
5.1.2.
It is known that, if $q$ is large enough, $Q_{T}^{G}(u)$ is a polynomial in $q$ with non-negative coefficients [Springer84, Theorem 5.2]. In fact, as shown in [GoodwinRohrle, Theorem 3.10], under the assumption that $Z$ is connected and $p$ is good for $G$, $Q_{T}^{G}(u)$ is a polynomial in $q$, unless $G$ has a factor of type $E_{8}$, in which case a quadratic base change may be necessary.
Convention 26.
We assume henceforth that $q$ is large enough so that $Q_{T}^{L}(\mathcal{O})$ is a polynomial in $q$ with non-negative coefficients for all standard Levi subgroups $L\subseteq G$ and all unipotent orbits in $\mathfrak{l}$.
5.1.3. Springer isomorphism
Since $\mathrm{char}(k)$ is very good for $G$, a theorem of Springer states that there exists a $G$-equivariant isomorphism $\varpi$ from the nilpotent cone of $\mathfrak{g}$ to the unipotent variety of $G$, cf. [Letellier05, §2.7.5].
5.1.4.
Let $\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}$ denote the space of $G^{F}$-invariant functions on $\mathfrak{g}^{F}$. For each $x\in\mathfrak{g}^{F}$, let $1_{x}^{G}$ denote the characteristic function of the adjoint orbit $G^{F}.x\subseteq\mathfrak{g}^{F}$. The set ${\{1_{x}^{G}\,|\,x\in\mathfrak{g}^{F}/G^{F}\}}$ is a basis of $\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}$. One can think of this set as the additive analogue of the set of irreducible complex characters of $G^{F}$. The Fourier transform $\mathcal{F}:\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}\rightarrow\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}$ is defined by
$\mathcal{F}(\phi)(x):=\sum_{y\in\mathfrak{g}^{F}}\psi(\kappa(x,y))\phi(y).$
5.1.5.
We write $x=x_{s}+x_{n}$ for the Jordan decomposition of $x\in\mathfrak{g}^{F}$. Define the Harish-Chandra induction map
${R_{\mathfrak{t}}^{\mathfrak{g}}:\mathbb{C}[\mathfrak{t}^{F}]\rightarrow\mathbb{C}[\mathfrak{g}^{F}]^{G^{F}}}$
by
$R_{\mathfrak{t}}^{\mathfrak{g}}(f)(x)=\frac{1}{|G_{x_{s}}^{F}|}\sum_{\{g\in G^{F}\,|\,g.x_{s}\in\mathfrak{t}^{F}\}}Q_{T}^{G_{x_{s}}}(\varpi(x_{n}))f(g.x_{s}).$
Here, $G_{x_{s}}$ is the centraliser of $x_{s}$ in $G$.
Definition 27.
We call $x\in\mathfrak{g}^{F}$ split if $x_{s}$ is $G^{F}$-conjugate to an element of $\mathfrak{t}^{F}$.
It follows immediately from the definition that if $R_{\mathfrak{t}}^{\mathfrak{g}}(f)(x)$ is non-zero, then $x$ is split.
5.2. A theorem of Kazhdan and Letellier
Let $f_{H}:\mathfrak{t}^{F}\rightarrow\mathbb{C}$ denote the character $\psi(\kappa|_{\mathfrak{t}}(-,H))$. The following result was proved by Kazhdan when $\mathrm{char}(k)$ is large. It was proved by Letellier that the characteristic being very good is sufficient, cf. [Letellier05, Theorem 7.3.3].
Theorem 28.
Suppose $H\in\mathfrak{t}^{F}$ is regular and $x\in\mathfrak{g}^{F}$. Then,
$\mathcal{F}(1_{H}^{G})(x)=q^{|\Phi^{+}|}R_{\mathfrak{t}}^{\mathfrak{g}}(f_{H})(x).$
Corollary 29.
If $H\in\mathfrak{t}^{F}$ is regular and $x\in\mathfrak{g}^{F}$, then $\mathcal{F}(1_{H}^{G})(x)\neq 0$ only if $x$ is split.
These results can be considered as additive analogues of Theorem 17 and Corollary 18.
5.2.1.
Let $(\mathfrak{g}^{F})^{\circ}$ denote the set of split elements of $\mathfrak{g}^{F}$. The above discussions implies
$|{A}^{F}|=|\mathfrak{g}^{F}|^{g-1}\sum_{x\in(\mathfrak{g}^{F})^{\circ}}|\mathfrak{g}_{x}^{F}|^{g}\mathcal{F}(1_{H}^{G})(x).$
5.2.2.
Let $x\in(\mathfrak{g}^{F})^{\circ}$ be a split element. Then $x_{s}$ is $G^{F}$-conjugate to an element $t\in\mathfrak{t}^{F}$, well-defined up to $W$-conjugation. So, the centraliser $G_{x_{s}}$ is a standard Levi subgroup of $G$, well-defined up to the action of $W$. Note that since $x_{n}$ and $x_{s}$ commute, we have $x_{n}\in\mathfrak{g}_{x_{s}}$; thus, we can consider the nilpotent orbit $G_{x_{s}}^{F}.x_{n}\subseteq\mathfrak{g}_{x_{s}}^{F}$.
5.2.3. Types
Consider the set of pairs $(L,\mathcal{N})$, where $L$ is a standard Levi subgroup of $G$ and $\mathcal{N}$ is the $L^{F}$-orbit of a nilpotent element of $\mathfrak{l}^{F}$, where $\mathfrak{l}=\mathrm{Lie}(L)$. Note that $W$ acts on this set by conjugation. Let $\mathcal{T}(\mathfrak{g})$ denote the set of $W$-orbits. We call the elements of $\mathcal{T}(\mathfrak{g})$ $\mathfrak{g}$-types and denote them by $\tau=[L,\mathcal{N}]$. By the above discussion, we have a map
$(\mathfrak{g}^{F})^{\circ}\rightarrow\mathcal{T}(\mathfrak{g}).$
This is the additive analogue of the map (4).
5.2.4.
The following proposition implies that $\mathcal{T}(\mathfrak{g})$ is a finite set and independent of the ground field:
Proposition 30.
There exists a finite set, independent of the finite field $\mathbb{F}_{q}$ (of good characteristic), parameterising nilpotent $L^{F}$-orbits in $\mathfrak{l}^{F}$.
Proof.
First of all, since $p$ is assumed to be good for $\mathfrak{g}$, it is also good for $\mathfrak{l}$. In this case, it is known that nilpotent orbits in $\mathfrak{l}$ are in bijection with nilpotent orbits in $\mathfrak{l}(\mathbb{C})$, cf. [Premet03]. Next, for any element $x\in\mathfrak{l}$, consider the orbit $L.x$. The Frobenius fixed point $(L.x)^{F}$ splits into $N_{x}$ many $L^{F}$-orbits, where $N_{x}:=|H^{1}(F,\pi_{0}(L_{x}))|$, cf. [DM20, Proposition 4.2.14]. Finally, observe that $N_{x}$ is the number of $F$-conjugacy classes in $\pi_{0}(L_{x})$. By Corollary 8, $p\nmid|\pi_{0}(L_{x})|$. Thus, the number of $F$-conjugacy classes in $\pi_{0}(L_{x})$ is independent of $k$. ∎
One can think of the above proposition as the additive analogue of the fact that unipotent representations of $G^{F}$ have a characteristic free description.
5.2.5.
Let $\tau=[L,\mathcal{N}]$ be a type. Recall that this means that $\mathcal{N}$ is the $L^{F}$-orbit of a nilpotent element $n\in\mathfrak{l}^{F}$. Choose $n\in\mathcal{N}$ and let $d(\mathcal{N}):=\dim(L.n)$. Clearly, this is independent of the choice of $n\in\mathcal{N}$. Let
$d(\tau):=\dim(L)-d(\mathcal{N}).$
The following is the additive analogue of Proposition 20:
Lemma 31.
Let $x\in(\mathfrak{g}^{F})^{\circ}$ be an element of type $\tau=[L,\mathcal{N}]$. Then
$\dim(\mathfrak{g}_{x})=d(\tau).$
Proof.
Observe that $\mathfrak{g}_{x}=\mathfrak{g}_{x_{s}}\cap\mathfrak{g}_{x_{n}}=C_{\mathfrak{g}_{x_{s}}}(x_{n})$. Indeed, if an element of $\mathfrak{g}$ commutes with $x$, then it must commute with $x_{s}$ and $x_{n}$, because these can be written as polynomials in $x$. Thus, we have
$\dim(\mathfrak{g}_{x})=\dim(C_{\mathfrak{g}_{x_{s}}}(x_{n}))=\dim(C_{G_{x_{s}}}(x_{n}))=\dim(G_{x_{s}})-\dim(G_{x_{s}}.{x_{n}})\\ =\dim(\mathfrak{g}_{x_{s}})-\dim(G_{x_{s}}.{x_{n}}).\qed$
5.2.6.
In view of the above discussions, we can rewrite the formula for $|{Y}^{F}|$ as follows
$|{Y}^{F}|=\frac{|Z^{F}||\mathfrak{g}^{F}|^{g-1}}{|G^{F}|}\sum_{\tau\in\mathcal{T}(\mathfrak{g})}q^{d(\tau)g}\sum_{x\in\mathfrak{g}^{F}_{\tau}}\mathcal{F}(1_{H}^{G})(x),$
where $\mathfrak{g}^{F}_{\tau}$ denotes the set of elements in $\mathfrak{g}^{F}$ of type $\tau$. For ease of notation, let
$\boxed{H_{\tau}(q):=\displaystyle\sum_{x\in\mathfrak{g}^{F}_{\tau}}\mathcal{F}(1_{H}^{G})(x).}$
Then we see that determining $|{Y}^{F}|$ is reduced to computing $H_{\tau}(q)$. Note that $H_{\tau}$ is the additive analogue of $S_{\tau}$ considered in §4.1.9.
5.3. Evaluating the sum $H_{\tau}$
First, observe that $\mathcal{F}(1_{H}^{G})$ is $G^{F}$-invariant; thus, instead of summing over all elements in $\mathfrak{g}^{F}_{\tau}$, we can sum over $G^{F}$-orbits $\mathfrak{O}_{x}\subseteq\mathfrak{g}^{F}$ of elements of $x\in\mathfrak{g}^{F}_{\tau}$. In other words,
$H_{\tau}(q)=\displaystyle\sum_{\mathfrak{O}\in\mathfrak{g}^{F}_{\tau}/G^{F}}|\mathfrak{O}|\mathcal{F}(1_{H}^{G})(\mathfrak{O}).$
We first show that the size of the orbit $|\mathfrak{O}|$ depends only on the type of $\mathfrak{O}$:
Lemma 32.
If $\mathfrak{O}$ has type $\tau=[L,\mathcal{N}]$, then $|\mathfrak{O}|=\frac{|G^{F}||\mathcal{N}|}{|L^{F}|}$.
Proof.
Let $x=x_{s}+x_{n}\in\mathfrak{g}^{F}$. Jordan decomposition is preserved under adjoint action; thus, for $g\in G$, $g.x=g.x_{s}+g.x_{n}$ is the Jordan decomposition of $g.x$. It follows that $g.x=x$ if and only if $g.x_{s}=x_{s}$ and $g.x_{n}=x_{n}$. In other words,
$G_{x}=G_{x_{s}}\cap G_{x_{n}}=C_{G_{x_{s}}}(x_{n}).$
Now if $x\in\mathfrak{g}^{F}$ has type $\tau=[L,\mathcal{N}]$, then $C_{G_{x_{s}}}(x_{n})=|L^{F}|/|\mathcal{N}|$. Thus, if $\mathfrak{O}$ is the $G^{F}$-orbit of $x$, we conclude
$|\mathfrak{O}|=\frac{|G^{F}|}{|G_{x}^{F}|}=\frac{|G^{F}|}{|C_{G_{x_{s}}}(x_{n})^{F}|}=\frac{|G^{F}||\mathcal{N}|}{|L^{F}|}.\qed$
The above lemma implies
$H_{\tau}(q)=\frac{|G^{F}||\mathcal{N}|}{|L^{F}|}\displaystyle\sum_{\mathfrak{O}\in\mathfrak{g}^{F}_{\tau}/G^{F}}\mathcal{F}(1_{H}^{G})(\mathfrak{O}).$
To proceed further, we need to better understand the values of the Fourier transform.
5.3.1.
Recall that Theorem 28 states $\mathcal{F}(1_{H}^{G})(x)=q^{|\Phi^{+}|}R_{\mathfrak{t}}^{\mathfrak{g}}(f_{H})(x)$ and that $R_{\mathfrak{t}}^{\mathfrak{g}}$ is defined via a sum over the set
$\mathcal{A}(x):=\{g\in G^{F}\,|\,g.x_{s}\in\mathfrak{t}^{F}\}.$
If we assume $x_{s}\in\mathfrak{t}^{F}$, then one readily checks
$\mathcal{A}(x)=\bigcup_{w\in W}\dot{w}G_{x_{s}}^{F}=\bigsqcup_{w\in W/W_{x_{s}}}\dot{w}G_{x_{s}}^{F}.$
Therefore, under this assumption we have:
$R_{\mathfrak{t}}^{\mathfrak{g}}(f_{H})(x)=\frac{Q_{T}^{G_{x_{s}}}(\varpi(x_{n}))}{|G_{x_{s}}^{F}|}\sum_{g\in\mathcal{A}(x)}f_{H}(g.x_{s})=\frac{Q_{T}^{G_{x_{s}}}(\varpi(x_{n}))}{|W_{x_{s}}|}\sum_{w\in W}f_{H}(w.x_{s}).$
5.3.2.
The above discussion implies that for a type $\tau=[L,\mathcal{N}]$, we have
$H_{\tau}=\frac{|G^{F}||\mathcal{N}|}{|L^{F}|}\sum_{x\in\mathfrak{g}^{F}_{\tau}/G^{F}}{q^{|\Phi^{+}|}}R_{\mathfrak{t}}^{\mathfrak{g}}(f_{H})(x)=\frac{|G^{F}||\mathcal{N}|}{|L^{F}|}\frac{q^{|\Phi^{+}|}Q_{T}^{L}(\varpi(\mathcal{N}))}{|W(L)|}\sum_{w\in W}\,\,\sum_{x\in\mathfrak{g}^{F}_{\tau}/G^{F}}f_{H}(w.x_{s}).$
Here, $x$ is an orbit representative satisfying $x_{s}\in\mathfrak{t}^{F}$.
5.3.3.
The element $x_{s}$ is well-defined up to $W/W_{x_{s}}$-conjugacy, thus, we can now rewrite the above sum as:
$\sum_{w\in W}\sum_{x\in\mathfrak{g}^{F}_{\tau}/G^{F}}f_{H}(w.x_{s})=\frac{|W(L)|}{|W|}\sum_{w\in W}\sum_{{\begin{subarray}{c}y\in\mathfrak{t}^{F}\\ [W_{y}]=[W(L)]\end{subarray}}}f_{H}(w.y)=\frac{|W(L)|}{|N_{W}(W(L))|}\sum_{w\in W}\sum_{{\begin{subarray}{c}y\in\mathfrak{t}^{F}\\ W_{y}=W(L)\end{subarray}}}f_{H}(w.y).$
5.3.4.
For ease of notation, let
$\beta_{L,H}(q):=\sum_{\begin{subarray}{c}y\in\mathfrak{t}^{F}\\ W_{y}=W(L)\end{subarray}}f_{H}(y).$
In view of the fact that $|[L]|=|W/N_{W}(W(L))|$ (see §4.2.3), we conclude:
$\boxed{H_{\tau}(q)=\frac{q^{|\Phi^{+}|}\,Q_{T}^{L}(\varpi(\mathcal{N}))\,|\mathcal{N}|\,|G^{F}|\,|[L]|}{|L^{F}|\,|W|}\sum_{w\in W}\beta_{L,w.H}(q).}$
Thus, the problem of computing $H_{\tau}$ reduces to evaluating $\beta_{L,H}$. Note that this is the additive analogue of $\alpha_{L,S}$ considered in the previous section.
5.4. Computing $\beta_{L,H}$
Let
$\beta_{L,H}^{\supseteq}:=\sum_{\begin{subarray}{c}t\in\mathfrak{t}^{F}\\ W_{t}\supseteq W(L)\end{subarray}}f_{H}(t)=\sum_{t\in(\mathfrak{t}^{F})^{W(L)}}f_{H}(t).$
Let $\mu$ denote the Möbius function on the poset of standard Levi subgroups of $G$ ordered by inclusion. [^9] Note that this poset is not the same as the one used in the previous section, but we have abused notation and denoted their Möbius functions by the same symbol. Möbius inversion gives
$\beta_{L,H}=\sum_{L^{\prime}\supseteq L}\mu(L,L^{\prime})\beta_{L^{\prime},H}^{\supseteq}$
5.4.1.
We are reduced to computing the Möbius function and the functions $\beta_{L,H}^{\supseteq}$. The latter can be computed as follows. Consider the canonical quotient map to the group of coinvariants
$f_{\mathfrak{l}}:\mathfrak{t}^{F}\rightarrow(\mathfrak{t}^{F})^{W(L)}=\mathfrak{t}^{F}/\mathfrak{t}^{F}\cap[\mathfrak{l}^{F},\mathfrak{l}^{F}].$
Then Pontryagin duality implies
$\beta_{L,H}^{\supseteq}=\begin{cases}|\mathfrak{t}^{F}/\mathfrak{t}^{F}\cap[\mathfrak{l}^{F},\mathfrak{l}^{F}]|&\textrm{if $f_{\mathfrak{l}}(H)=0$; i.e., $H\in[\mathfrak{l}^{F},\mathfrak{l}^{F}]$;}\\ 0&\textrm{otherwise}.\end{cases}$
5.4.2.
Since $H$ is generic, we have that $H\in[\mathfrak{l},\mathfrak{l}]$ if and only if $\mathfrak{l}=\mathfrak{g}$. Note that
$|\mathfrak{t}^{F}/\mathfrak{t}^{F}\cap[\mathfrak{g}^{F},\mathfrak{g}^{F}]|=|Z(\mathfrak{g})^{F}|.$
Thus, $\beta_{L,H}^{\supseteq}$ is non-zero only if $L=G$, in which case $\mathfrak{t}^{F}/(\mathfrak{t}^{F}\cap[\mathfrak{l}^{F},\mathfrak{l}^{F}])=Z(\mathfrak{g})^{F}=\mathrm{Lie}(Z)^{F}$. We conclude
$\boxed{\beta_{L,H}=q^{\dim(Z)}\mu(L,G).}$
This is the additive analogue of (5). Observe that $\beta_{L,H}$ is independent of $H$ (provided it is generic).
5.5. Conclusion of the point count for once-punctured case
We have seen that
$|{Y}^{F}|=\frac{|Z^{F}||\mathfrak{g}^{F}|^{g-1}}{|G^{F}|}\sum_{\tau\in\mathcal{T}(\mathfrak{g})}q^{gd(\tau)}H_{\tau}(q),$
where
$H_{\tau}(q)=q^{|\Phi^{+}|+\dim(Z)}Q_{T}^{L}(\varpi(\mathcal{N}))|\mathcal{N}|\mu(L,G)|[L]|\frac{|G^{F}|}{|L^{F}|}.$
5.5.1.
It is clear from this description that $|{Y}^{F}|$ is a rational function in $q$. If we replace the Frobenius $F$ by $F^{m}$ for some positive integer $m$, then $|{Y}^{F^{m}}|$ is given by the same rational function with $q$ replaced by $q^{m}$. It follows that this rational function is actually a polynomial. Thus, ${Y}$ is polynomial count, in the sense of [LetellierVillegas], and our expression for $|{Y}^{F}|$ is the $E$-polynomial of $Y$. The fact that it is monic is left as an exercise to the reader; see [Giannini24] for details.
5.6. Multi-punctured case
Let ${Y}$ be the additive character variety associated to an $n$-punctured surface group and $(H_{1},\ldots,H_{n})\in\mathfrak{t}^{n}$ where each $H_{i}$ is regular and the tuple is generic. Then a similar analysis as in the once-punctured case gives
$|{Y}^{F}|=\frac{|Z^{F}||\mathfrak{g}^{F}|^{g-1}}{|G^{F}|}\sum_{\tau\in\mathcal{T}(\mathfrak{g})}q^{gd(\tau)}H_{\tau}(q),$
where
$H_{\tau}(q)=q^{n|\Phi^{+}|+\dim(Z)}\frac{|G^{F}|}{|L^{F}|}|\mathcal{N}|\Big{(}Q_{T}^{L}(\varpi(\mathcal{N}))\Big{)}^{n}\bigg{(}\frac{|W|}{|W(L)|}\bigg{)}^{n-1}|[L]|\,\mu(L,G).$
5.6.1.
For the benefit of the reader, we recall what the notation means. Here, $Z$ is the centre of $G$, $\mathcal{T}(\mathfrak{g})$ is the set of $\mathfrak{g}$-types (i.e., $W$-orbits of pairs $(L,\mathcal{N})$, where $L$ is a standard Levi subgroup of $G$ and $\mathcal{N}$ a nilpotent $L^{F}$-orbit in $\mathfrak{l}^{F}$), and $d(\tau)=\dim(L)-\dim(L.n)$ where $n\in\mathcal{N}$, $Q_{T}^{L}$ is the Green function evaluated at the unipotent class $\varpi(\mathcal{N})\subseteq L^{F}$ associated to $\mathcal{N}$, $|[L]|$ is the size of the $W$-orbit of $L$, and $\mu(L,G)$ is Möbius function associated to the poset of standard Levi subgroups of $G$. The same argument as above implies that $Y$ is polynomial count with the monic counting polynomial given by $|Y^{F}|$; see [Giannini24] for details.
6. Examples
In this section, we give examples of Euler characteristics and counting polynomials of the character variety $X$ and its additive analogue $Y$. We also give tables containing $G$-types, $\mathfrak{g}$-types, and the associated data required to compute counting polynomials. In the case $(g,n)=(0,3)$, we also give counting polynomials when $G$ is a group of high rank such as $F_{4}$ and $E_{6}$ (other choices for $(g,n)$ are computable using the code found at [Bailey]).
Throughout, $\Phi_{i}$ is the $i$th cyclotomic polynomial. In particular,
$\Phi_{1}=q-1,\quad\Phi_{2}=q+1,\quad\Phi_{3}=q^{2}+q+1,\quad\Phi_{4}=q^{2}+1,\quad\Phi_{6}=q^{2}-q+1.$
We use the notation of [Carter93, §13] for characters of Weyl groups, unipotent characters, and nilpotent orbits. We label Levi subgroups and endoscopy groups by their root systems, with $A_{1}^{\prime}$ denoting short copies of $A_{1}$. Given an isolated standard endoscopy group $L$ of $G$, we denote by $\pi_{0}^{L}$ the number of components $|\pi_{0}(Z(\check{L}))|$.
6.1. Euler characteristic of ${X}$
In this subsection, we give examples of Euler characteristics of the character variety $X$.
6.1.1. $g=1$ and trivial centre
Using Theorem 2, it is straightforward to show
$\chi({X})=|W|^{n-1}\sum_{L}|W(L)|\,|\mathrm{Irr}(W(L))|\,\nu(L),$
where the sum is over all endoscopy groups of $G$ containing $T$.
6.1.2.
Below we give tables containing the data required for computing this Euler characteristic in the cases $G=\mathrm{SO}_{5}$ and $G=G_{2}$. Using these tables, we find
$\chi({X}_{\mathrm{SO}_{5}})=72\times 8^{n-1}\qquad\text{and}\qquad\chi({X}_{G_{2}})=96\times 12^{n-1}.$
$[L]$ $|[L]|$ $W(L)$ $|W(L)|$ $|\mathrm{Irr}(W(L))|$ $\nu(L)$
$[\mathrm{SO}_{5}]$ $1$ $D_{8}$ $8$ $5$ $2$
$[{A_{1}\times A_{1}}]$ $1$ $S_{2}\times S_{2}$ $4$ $4$ $2$
$[{A_{1}}]$ $2$ $S_{2}$ $2$ $2$ $-4$
$[{A_{1}^{\prime}}]$ $2$ $S_{2}$ $2$ $2$ $-2$
$[T]$ $1$ $1$ $1$ $1$ $8$
Figure 2. Required data for $\chi(X)$ when $G=\mathrm{SO}_{5}$ and $g=1$.
$[L]$ $|[L]|$ $W(L)$ $|W(L)|$ $|\mathrm{Irr}(W(L))|$ $\nu(L)$
$[G_{2}]$ $1$ $D_{12}$ $12$ $6$ $1$
$[{A_{2}}]$ $1$ $S_{3}$ $6$ $3$ $2$
$[{A_{1}\times A_{1}^{\prime}}]$ $3$ $S_{2}\times S_{2}$ $4$ $4$ $1$
$[{A_{1}}]$ $3$ $S_{2}$ $2$ $2$ $-4$
$[{A_{1}^{\prime}}]$ $3$ $S_{2}$ $2$ $2$ $-2$
$[T]$ $1$ $1$ $1$ $1$ $12$
Figure 3. Required data for $\chi(X)$ when $G=G_{2}$ and $g=1$.
6.1.3. $g=0$
Using Theorem 2, it is straightforward to show the Euler characteristic is
$\chi({X})=\frac{1}{(2r)!}\frac{d^{2r}}{dq^{2r}}\bigg{|}_{q=1}\xi(q),$
where $r:=\dim(T)-\dim(Z)$ is the semisimple rank of $G$ and
$\xi(q):=q^{|\Phi(G)^{+}|(n-2)}\sum_{L}\bigg{(}\frac{|W|}{|W(L)|}\bigg{)}^{n-1}\nu(L)\sum_{\rho}\dim(\rho)^{n}\bigg{(}\frac{P_{W(L)}(q)}{\dim(\tilde{\rho})}\bigg{)}^{n-2}.$
Above, the first sum is over all endoscopy groups of $G$ containing $T$, and the second sum is over all irreducible characters of $W(L)$. We give a few examples of Euler characteristics below. Note that the Euler characteristic of ${X}$ is 00 if $n=1,2$, which is consistent with the fact that ${X}$ is empty in these cases.
$G$ $\chi({X})$
$\mathrm{GL}_{2}$ $2^{n-4}(n-1)(n-2)$
$\mathrm{GL}_{3}$ $2^{n-5}3^{n-3}(n-1)(n-2)(9n^{2}-27n+16)$
$\mathrm{GL}_{4}$ $2^{3n-9}3^{n-4}(n-1)(n-2)(108n^{4}-648n^{3}+1350n^{2}-1129n+324)$
$\mathrm{SO}_{5}$ $2^{3n-8}(n-1)(n-2)(11n^{2}-33n+19)$
$G_{2}$ $2^{2n-7}3^{n-3}(n-1)(n-2)(207n^{2}-621n+350)$
Figure 4. Euler characteristics of ${X}$ when $g=0$.
6.2. Euler characteristic of ${Y}$
We now consider the additive character variety of Theorem 5. Using the counting polynomial provided by the theorem, it is straightforward to show
$\chi({Y})=\frac{1}{r!}\frac{d^{r}}{dq^{r}}\bigg{|}_{q=1}\zeta(q),$
where $r:=\dim(T)-\dim(Z)$ is the semisimple rank of $G$ and
$\zeta(q):=\frac{|\mathfrak{g}(k)|^{g-1}}{q^{|\Phi(G)^{+}|}P_{W}(q)}\sum_{\tau\in\mathcal{T}(\mathfrak{g})}q^{gd(\tau)}H_{\tau}(q).$
6.2.1.
We give a few examples of Euler characteristics below. As before, the Euler characteristic of ${Y}$ is 0 if $(g,n)=(0,1)$ or $(0,2)$ since $Y$ is empty in these cases.
$G$ $\chi({Y})$
$\mathrm{GL}_{2}$ $1+2^{n-2}(4g+n-3)$
$\mathrm{GL}_{3}$ $1+2\cdot 3^{n-1}(3g+n-3)$ $+3^{n-2}2^{n+3}(144g^{2}+72g(n-3)+80-53n+9n^{2})$
$\mathrm{SO}_{5}$ $1+2^{n-2}+2^{2n-3}(24g+7n-22)$ $+2^{3n-5}(96g^{2}+48g(n-3)+53-35n+6n^{2})$
$G_{2}$ $\frac{1}{6}(6-2^{n}+19\cdot 3^{n+1}+5^{n})$ $+395\cdot 12^{n-2}+5g\cdot 6^{n}+6^{n-2}(48-65\cdot 2^{n})n$ $+5\cdot 3^{n}4^{n-2}(16g^{2}+8g(n-3)+n^{2})$
Figure 5. Euler characteristics of ${Y}$.
6.3. The group $G=\mathrm{SO}_{5}$
$\tau=[L,\rho]$ $m_{\tau}(q)$ $S_{\tau}(q)$ $|L^{F}|$ $\tilde{\rho}(1)$ $\rho(1)$ $|W(L)|$ $|[L]|$ $\pi_{0}^{L}$ $\nu(L)$
$[\mathrm{SO}_{5},\binom{2}{}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $2\Phi_{1}^{2}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $1$ $1$ $8$ $1$ $2$ $2$
[…]
$[{A_{1}\times A_{1}},2^{1}\otimes 2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $1$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}\times A_{1}},2^{1}\otimes 1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}\times A_{1}},1^{2}\otimes 2^{1}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}\times A_{1}},1^{2}\otimes 1^{2}]$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $2^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q^{2}$ $1$ $4$ $1$ $4$ $2$
$[{A_{1}},2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 4^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $2$ $2$ $-4$
$[{A_{1}},1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 4^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $2$ $2$ $-4$
$[{A_{1}^{\prime}},2^{1}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}$ $-4^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $2$ $1$ $-2$
$[{A_{1}^{\prime}},1^{2}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}$ $-4^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $2$ $1$ $-2$
$[T,1^{1}]$ $q^{4}\Phi_{1}^{2}$ $8^{n}\Phi_{1}^{2}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $1$ $8$
Figure 6. The fourteen $\mathrm{SO}_{5}$-types.
$\tau=[L,\mathcal{N}]$ $q^{d(\tau)}$ $H_{\tau}(q)$ $|L^{F}|$ $|\mathcal{N}|$ $Q_{T}^{L}(\varpi(\mathcal{N}))$ $|W(L)|$ $|[L]|$ $\mu(L,G)$
$[\mathrm{SO}_{5},1^{5}]$ $q^{10}$ $q^{4n}\Phi_{2}^{2n}\Phi_{4}^{n}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $1$ $\Phi_{2}^{2}\Phi_{4}$ $8$ $1$ $1$
$[\mathrm{SO}_{5},2^{2}1^{1}]$ $q^{6}$ $q^{4n}\Phi_{1}\Phi_{2}^{2n+1}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\Phi_{1}\Phi_{2}\Phi_{4}$ $\Phi_{2}^{2}$ $8$ $1$ $1$
$[\mathrm{SO}_{5},3^{1}1^{2}]$ $q^{4}$ $\frac{1}{2}q^{4n+1}\Phi_{1}\Phi_{2}^{2}\Phi_{4}(3q+1)^{n}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\frac{1}{2}q\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $13q+$ 1 $8$ $1$ $1$
$[\mathrm{SO}_{5},3^{1}1^{2}_{\star}]$ $q^{4}$ $\frac{1}{2}q^{4n+1}\Phi_{1}^{2}\Phi_{2}^{n+1}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $\frac{1}{2}q\Phi_{1}^{2}\Phi_{2}\Phi_{4}$ $\Phi_{2}$ $8$ $1$ $1$
$[\mathrm{SO}_{5},5^{1}]$ $q^{2}$ $q^{4n+2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{4}$ $1$ $8$ $1$ $1$
$[A_{1},1^{2}]$ $q^{4}$ $-2\cdot 4^{n-1}q^{4n+3}\Phi_{2}^{n+1}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $\Phi_{2}$ $2$ $2$ $-1$
$[A_{1},2^{1}]$ $q^{2}$ $-2\cdot 4^{n-1}q^{4n+3}\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ $1$ $2$ $2$ $-1$
$[A_{1}^{\prime},1^{2}]$ $q^{4}$ $-2\cdot 4^{n-1}q^{4n+3}\Phi_{2}^{n+1}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $\Phi_{2}$ $2$ $2$ $-1$
$[A_{1}^{\prime},2^{1}]$ $q^{2}$ $-2\cdot 4^{n-1}q^{4n+3}\Phi_{1}\Phi_{2}^{2}\Phi_{4}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ $1$ $2$ $2$ $-1$
$[T,0]$ $q^{2}$ $3\cdot 8^{n-1}q^{4n+4}\Phi_{2}^{2}\Phi_{4}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $3$
Figure 7. The ten $\mathfrak{so}_{5}$-types.
$(g,n)$ $|{X}(\mathbb{F}_{q})|$
$(0,3)$ $22q^{4}+12q^{3}+48q^{2}+12q+$ 2
$(0,4)$ $2q^{12}+16q^{11}+68q^{10}+208q^{9}+530q^{8}$ $+1216q^{7}+1968q^{6}+\cdots+16q+2$
$(1,1)$ $2q^{8}+4q^{7}+6q^{6}+16q^{5}+16q^{4}$ $+16q^{3}+6q^{2}+4q+2$
$(1,2)$ $2q^{16}+8q^{15}+16q^{14}+24q^{13}+32q^{12}+40q^{11}$ $+48q^{10}+24q^{9}+188q^{8}+\cdots+8q+2$
$(2,1)$ $2q^{28}+4q^{27}-4q^{25}-8q^{24}-12q^{23}+2q^{22}$ $+16q^{21}+6q^{20}+52q^{19}-40q^{18}+228q^{17}$ $-840q^{16}+708q^{15}-228q^{14}+\cdots+4q+2$
Figure 8. Examples of the polynomial $|{X}(\mathbb{F}_{q})|$ when $G=\mathrm{SO}_{5}$.
$(g,n)$ $|{Y}(\mathbb{F}_{q})|$
$(0,3)$ $q^{4}+6q^{3}+20q^{2}$
$(0,4)$ $q^{12}+8q^{11}+33q^{10}+96q^{9}+223q^{8}+440q^{7}+548q^{6}$
$(1,1)$ $q^{8}+2q^{7}+4q^{6}+4q^{5}+q^{4}$
$(1,2)$ $q^{16}+4q^{15}+9q^{14}+16q^{13}+25q^{12}$ $+36q^{11}+36q^{10}+16q^{9}+q^{8}$
$(2,1)$ $q^{28}+2q^{27}+3q^{26}+4q^{25}+5q^{24}+6q^{23}+8q^{22}$ $+10q^{21}+11q^{20}+10q^{19}+7q^{18}+4q^{17}+q^{16}$
Figure 9. Examples of the polynomial $|{Y}(\mathbb{F}_{q})|$ when $\mathfrak{g}=\mathfrak{so}_{5}$.
6.4. The group $G=G_{2}$
$\tau=[L,\rho]$ $m_{\tau}(q)$ $S_{\tau}(q)$ $|L^{F}|$ $\rho(1)$ $\tilde{\rho}(1)$ $|W(L)|$ $|[L]|$ $\pi_{0}^{L}$ $\nu(L)$
$[G_{2},\phi_{1,0}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{1}^{2}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $1$ $1$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{1,3}^{\prime}]$ $3q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $\Phi_{1}^{2}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q\Phi_{3}\Phi_{6}$ $1$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{1,3}^{\prime\prime}]$ $3q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $\Phi_{1}^{2}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q\Phi_{3}\Phi_{6}$ $1$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{2,1}]$ $6q^{5}\Phi_{1}^{2}\Phi_{6}$ $2^{n}\Phi_{1}^{2}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{6}q\Phi_{2}^{2}\Phi_{3}$ $2$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{2,2}]$ $2q^{5}\Phi_{1}^{2}\Phi_{3}$ $2^{n}\Phi_{1}^{2}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{2}q\Phi_{2}^{2}\Phi_{6}$ $2$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{1,6}]$ $\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{1}^{2}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{6}$ $1$ $12$ $1$ $1$ $1$
$[{A_{2}},3^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $2^{n}\Phi_{1}^{2}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $1$ $1$ $6$ $1$ $3$ $2$
$[{A_{2}},2^{1}1^{1}]$ $q^{5}\Phi_{1}^{2}\Phi_{3}$ $4^{n}\Phi_{1}^{2}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $q\Phi_{2}$ $2$ $6$ $1$ $3$ $2$
$[{A_{2}},1^{3}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $2^{n}\Phi_{1}^{2}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $q^{3}$ $1$ $6$ $1$ $3$ $2$
$[{A_{1}\times A_{1}^{\prime}},2^{1}\otimes 2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $1$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}\times A_{1}^{\prime}},2^{1}\otimes 1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}\times A_{1}^{\prime}},1^{2}\otimes 2^{1}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}\times A_{1}^{\prime}},1^{2}\otimes 1^{2}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q^{2}$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}},2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 6^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $3$ $1$ $-4$
$[{A_{1}},1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 6^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $3$ $1$ $-4$
$[{A_{1}^{\prime}},2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}$ $-6^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $3$ $1$ $-2$
$[{A_{1}^{\prime}},1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}$ $-6^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $3$ $1$ $-2$
$[T,1^{1}]$ $q^{6}\Phi_{1}^{2}$ $12^{n}\Phi_{1}^{2}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $1$ $12$
Figure 10. The eighteen ${G}_{2}$-types.
$\tau=[L,\mathcal{N}]$ $q^{d(\tau)}$ $H_{\tau}(q)$ $|L^{F}|$ $|\mathcal{N}|$ $Q_{T}^{L}(\varpi(\mathcal{N}))$ $|W(L)|$ $|[L]|$ $\mu(L,G)$
$[G_{2},0]$ $q^{14}$ $q^{6n}\Phi_{2}^{2n}\Phi_{3}^{n}\Phi_{6}^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $1$ $\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $12$ $1$ $1$
$[G_{2},A_{1}]$ $q^{8}$ $q^{6n}\Phi_{1}\Phi_{2}^{n+1}\Phi_{3}^{n+1}\Phi_{6}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{1}\Phi_{2}\Phi_{3}\Phi_{6}$ $\Phi_{2}\Phi_{3}$ $12$ $1$ $1$
$[G_{2},\tilde{A_{1}}]$ $q^{6}$ $q^{6n+2}\Phi_{1}\Phi_{2}^{n+1}\Phi_{3}\Phi_{6}(2q+1)^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{2}\Phi_{1}\Phi_{2}\Phi_{3}\Phi_{6}$ $(2q+1)\Phi_{2}$ $12$ $1$ $1$
$[G_{2},G_{2}(a_{1})_{1^{3}}]$ $q^{4}$ $\frac{1}{6}q^{6n+2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}(4q+1)^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{6}q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $14q+$ 1 $12$ $1$ $1$
$[G_{2},G_{2}(a_{1})_{2^{1}1^{1}}]$ $q^{4}$ $\frac{1}{2}q^{6n+2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}(2q+1)^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{2}q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $12q+$ 1 $12$ $1$ $1$
$[G_{2},G_{2}(a_{1})_{3^{1}}]$ $q^{4}$ $\frac{1}{3}q^{6n+2}\Phi_{1}^{2}\Phi_{2}^{n+2}\Phi_{3}\Phi_{6}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q^{2}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{2}$ $12$ $1$ $1$
$[G_{2},G_{2}]$ $q^{2}$ $q^{6n+4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $1$ $12$ $1$ $1$
$[A_{1},1^{2}]$ $q^{4}$ $-3\cdot 6^{n-1}q^{6n+5}\Phi_{2}^{n+1}\Phi_{3}\Phi_{6}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $\Phi_{2}$ $2$ $3$ $-1$
$[A_{1},2^{1}]$ $q^{2}$ $-3\cdot 6^{n-1}q^{6n+5}\Phi_{1}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ $1$ $2$ $3$ $-1$
$[A_{1}^{\prime},1^{2}]$ $q^{4}$ $-3\cdot 6^{n-1}q^{6n+5}\Phi_{2}^{n+1}\Phi_{3}\Phi_{6}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $\Phi_{2}$ $2$ $3$ $-1$
$[A_{1}^{\prime},2^{1}]$ $q^{2}$ $-3\cdot 6^{n-1}q^{6n+5}\Phi_{1}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q\Phi_{1}^{2}\Phi_{2}$ $\Phi_{1}\Phi_{2}$ $1$ $2$ $3$ $-1$
$[T,0]$ $q^{2}$ $5\cdot 12^{n-1}q^{6n+6}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $5$
Figure 11. The twelve $\mathfrak{g}_{2}$-types.
$(g,n)$ $|{X}(\mathbb{F}_{q})|$
$(0,3)$ $q^{8}+6q^{7}+20q^{6}+58q^{5}+180q^{4}+58q^{3}+20q^{2}+6q+1$
$(0,4)$ $q^{20}+8q^{19}+34q^{18}+104q^{17}+259q^{16}+560q^{15}+1106q^{14}$ $+2080q^{13}+3890q^{12}+7440q^{11}+11444q^{10}+\cdots+8q+1$
$(1,1)$ $q^{12}+2q^{11}+2q^{10}+4q^{9}+9q^{8}+26q^{7}+8q^{6}+\cdots+2q+1$
$(1,2)$ $q^{24}+4q^{23}+8q^{22}+12q^{21}+16q^{20}+20q^{19}+26q^{18}+36q^{17}$ $+57q^{16}+88q^{15}+198q^{14}-240q^{13}+700q^{12}+\cdots+4q+1$
$(2,1)$ $q^{40}+2q^{39}-2q^{37}-q^{36}-3q^{34}-6q^{33}+8q^{31}+7q^{30}-4q^{28}$ $-2q^{27}-9q^{26}+482q^{25}-2885q^{24}+10278q^{23}-23499q^{22}$ $+35928q^{21}-40590q^{20}+\cdots+2q+1$
Figure 12. Examples of the polynomial $|{X}(\mathbb{F}_{q})|$ when $G=G_{2}$.
$(g,n)$ $|{Y}(\mathbb{F}_{q})|$
$(0,3)$ $q^{8}+6q^{7}+19q^{6}+45q^{5}+99q^{4}$
$(0,4)$ $q^{20}+8q^{19}+33q^{18}+96q^{17}+225q^{16}+456q^{15}$ $+831q^{14}+1392q^{13}+2191q^{12}+3300q^{11}+3498q^{10}$
$(1,1)$ $q^{12}+2q^{11}+3q^{10}+5q^{9}+9q^{8}+8q^{7}+2q^{6}$
$(1,2)$ $q^{24}+4q^{23}+9q^{22}+16q^{21}+25q^{20}+36q^{19}+49q^{18}$ $+64q^{17}+85q^{16}+110q^{15}+99q^{14}+40q^{13}+2q^{12}$
$(2,1)$ $q^{40}+2q^{39}+3q^{38}+4q^{37}+5q^{36}+6q^{35}+7q^{34}+8q^{33}+9q^{32}+11q^{31}$ $+13q^{30}+15q^{29}+18q^{28}+20q^{27}+21q^{26}+18q^{25}+12q^{24}+6q^{23}+q^{22}$
Figure 13. Examples of the polynomial $|{Y}(\mathbb{F}_{q})|$ when $\mathfrak{g}=\mathfrak{g}_{2}$.
6.5. Further examples
In this subsection, we give a few examples of counting polynomials obtained using our formula and the computer code [Bailey].
$G$ $|X(\mathbb{F}_{q})|$
$B_{3}$ $22q^{12}+18q^{11}+88q^{10}+312q^{9}+952q^{8}+2478q^{7}+4980q^{6}+\cdots+18q+$ 2
$C_{3}$ $22q^{12}+18q^{11}+88q^{10}+320q^{9}+970q^{8}+2506q^{7}+5060q^{6}+\cdots+18q+$ 2
$D_{4}$ $4q^{16}+48q^{15}+308q^{14}+1408q^{13}+5140q^{12}+16176q^{11}$ $+43748q^{10}+96256q^{9}+152864q^{8}+\cdots+48q+4$
$F_{4}$ $q^{40}+12q^{39}+77q^{38}+352q^{37}+1287q^{36}+4004q^{35}+11010q^{34}+27444q^{33}$ $+63155q^{32}+136096q^{31}+277849q^{30}+542780q^{29}+1023705q^{28}+1879344q^{27}$ $+3384741q^{26}+6009812q^{25}+10498044q^{24}+17873468q^{23}+29224321q^{22}$ $+44790488q^{21}+58508548q^{20}+\cdots+12q+1$
$E_{6}$ $3q^{60}+54q^{59}+510q^{58}+3366q^{57}+17442q^{56}+75579q^{55}+284829q^{54}+958746q^{53}$ $+2938005q^{52}+8314473q^{51}+21970542q^{50}+54685197q^{49}+129125661q^{48}+290973411q^{47}$ $+628951404q^{46}+1310014803q^{45}+2640078312q^{44}+5167269447q^{43}+9854714367q^{42}+18363440685q^{41}$ $+33501157887q^{40}+59902887102q^{39}+104998780206q^{38}+180252368460q^{37}$ $+302421156681q^{36}+493955088993q^{35}+780185905473q^{34}+1178431874241q^{33}$ $+1671208058577q^{32}+2158923553011q^{31}+2408708043594q^{30}+\cdots+510q^{2}+54q+3$
Figure 14. Formulas for $|X(\mathbb{F}_{q})|$ when $(g,n)=(0,3)$.
$G$ $|Y(\mathbb{F}_{q})|$
$B_{3}$ $q^{12}+9q^{11}+43q^{10}+147q^{9}+412q^{8}+948q^{7}+1535q^{6}$
$C_{3}$ $q^{12}+9q^{11}+43q^{10}+148q^{9}+413q^{8}+946q^{7}+1533q^{6}$
$D_{4}$ $q^{16}+12q^{15}+76q^{14}+340q^{13}+1206q^{12}+3668q^{11}+9451q^{10}+19080q^{9}+24492q^{8}$
$F_{4}$ $q^{40}+12q^{39}+76q^{38}+340q^{37}+1210q^{36}+3652q^{35}+9722q^{34}+23428q^{33}+52049q^{32}$ $+108072q^{31}+211964q^{30}+396184q^{29}+711295q^{28}+1235704q^{27}+2091762q^{26}$ $+3459608q^{25}+5550351q^{24}+8475196q^{23}+11932547q^{22}+14602160q^{21}+12631801q^{20}$
$E_{6}$ $q^{60}+18q^{59}+169q^{58}+1104q^{57}+5644q^{56}+24070q^{55}+89110q^{54}+294202q^{53}$ $+883118q^{52}+2445142q^{51}+6314277q^{50}+15342436q^{49}+35328016q^{48}+77552468q^{47}$ $+163141890q^{46}+330385400q^{45}+646787466q^{44}+1228563901q^{43}+2271409786q^{42}$ $+4097165169q^{41}+7220215707q^{40}+12432586601q^{39}+20893248670q^{38}$ $+34175389336q^{37}+54155977909q^{36}+82488035626q^{35}+119164974059q^{34}$ $+159544431246q^{33}+189727222032q^{32}+183584161672q^{31}+112360075923q^{30}$
Figure 15. Formulas for $|Y(\mathbb{F}_{q})|$ when $(g,n)=(0,3)$.